### v4

# Higher-weight Jacobians

Abstract.
We define and study Jacobians of Hodge structures with weight greater than 1. Jacobians of weight 2 naturally come up in the context of the Brauer group and the Tate conjecture. They were previously studied in a special case by Beauville in his work on surfaces of maximal Picard number, and are related to the work of Totaro on Hodge structures with no middle pieces. Higher-weight Jacobians are complex tori, and it is generally quite difficult to tell if they are algebraic. After discussing some general theory, we compute numerous examples of Jacobians of various weights for special classes of varieties: abelian varieties of maximal Picard number, Kummer varieties, and singular K3 surfaces. It turns out that all of these Jacobians are algebraic. We compute their fields of definition.
1. Introduction
Classically, one can realize the Jacobian of a complex curve $C$ as a complex torus using the exponential sequence:
${\rm H}^{0}(C,\mathscr{O})\to{\rm H}^{0}(C,\mathbf{G}_{m})\to{\rm H}^{1}(C,\mathbf{Z}(1))\to{\rm H}^{1}(C,\mathscr{O})\to{\rm H}^{1}(C,\mathbf{G}_{m})\to{\rm H}^{2}(C,\mathbf{Z}(1)).$
We know that the complex span of the image of ${\rm H}^{1}(C,\mathbf{Z}(1))$ is all of ${\rm H}^{1}(C,\mathscr{O})$, and we then get a complex torus using the fact that the image is discrete, which is immediate from the surjectivity of the exponential map $\mathbf{C}\to\mathbf{C}^{\times}$.
These elementary observations point in various directions. First, the cohomology has a connected part that is itself represented by a group scheme. Second, the integral cohomology sits discretely in the coherent cohomology, uniformizing this group scheme. Third, the full abelian variety thus constructed can be interpreted, in geometric terms, as parametrizing invertible sheaves on $C$. If we carry out the same analysis for a smooth projective complex variety $X$ of arbitrary dimension, we get the Albanese variety of $X$: the complex span of the integral cohomology is always the full coherent cohomology (by Hodge theory), and discreteness again follows from surjectivity of the exponential map. The deepest subtlety in this classical setting is the algebraicity of the resulting complex torus. In one version of the Albanese story, this ultimately relies on the fact that the Hodge structure has odd weight.
The question then arises in higher cohomological degrees: when can one attach an abelian variety to $X$ that measures the connected part of a cohomology group? To be specific, what happens for ${\rm H}^{2}$? We are immediately met by strange roadblocks: for smooth $X$, ${\rm H}^{2}(X,\mathbf{G}_{m})$ (in fact, all higher degree cohomology) is always torsion, equal to the Brauer group $\operatorname{Br}(X)$. Moreover, the map ${\rm H}^{2}(X,\mathbf{Z}(1))\to{\rm H}^{2}(X,\mathscr{O})$ need not have discrete image; that is only the case when the Picard number of $X$ is maximal. Finally, the cohomology has even degree, so there is no reason for any torus arising in this construction to be algebraic.
Notation 1.1.
Given a (pure) integral Hodge structure $H$ of weight $m$ such that $H^{i,j}=0$ for $i<0$ or $j<0$, define $\mathfrak{L}(H)$ to be the image of the natural map $H\to H^{0,m}$ and let $\mathfrak{C}(H)=H^{0,m}/\mathfrak{L}(H)$. If $H={\rm H}^{m}(X,\mathbf{Z})$ for a smooth complex projective variety $X$, we will write $\mathfrak{L}_{m}(X)=\mathfrak{L}(H)$ and $\mathfrak{C}_{m}(X)=\mathfrak{C}(H)$. In this case,
$\mathfrak{C}_{m}(X)={\rm H}^{m}(X,\mathscr{O}_{X})/\mathfrak{L}_{m}(X)={\rm coker}(\iota_{m}),$
where $\iota_{m}:{\rm H}^{m}(X,\mathbf{Z})\to{\rm H}^{m}(X,\mathscr{O}_{X})$ is the natural map induced by functoriality.
Definition 1.2.
We see that a pure integral Hodge structure $H$ of weight $m>0$ has a Jacobian if $\mathfrak{L}(H)$ is a discrete subgroup of the complex vector space $H^{0,m}$. If $H={\rm H}^{m}(X,\mathbf{Z})$ with $X$ a complex variety, we instead say that $X$ has a Jacobian of weight $m$, or an $m$-Jacobian. In this case, $\mathfrak{C}_{m}(X)$ is naturally a complex torus. When $m=2$, we will also use the name Brauer-Jacobian, and we will write $\mathfrak{C}$ for $\mathfrak{C}_{2}$.
In particular, when $C$ is a curve, the “weight $1$ Jacobian of $C$” defined above coincides with the realization of the Jacobian of $C$ as a complex torus via the exponential sequence. Furthermore, note that there is a natural inclusion
$\ker({\rm H}^{m}(X,\mathbf{G}_{m})\to{\rm H}^{m+1}(X,\mathbf{Z}(1)))\hookrightarrow\mathfrak{C}_{m}(X)$
that identifies the former with the torsion subgroup of the latter. That is, the topologically trivial degree $m$ cohomology is identified with the torsion points of the $m$-Jacobian; in particular, when $X$ has a $2$-Jacobian or Brauer-Jacobian, the torsion points of the Brauer-Jacobian are identified with the Brauer group of $X$ (whence the name).
In this paper, we study examples of Brauer-Jacobians for several types of varieties, with a focus on CM abelian varieties, where we can make the calculations quite explicit. This general phenomenon has been studied before, notably in a paper of Beauville [beauville], where various surfaces of maximal Picard number were discussed. There is also the fascinating work of Kollár and Joglar-Prieto [kollar2003] in the context of real algebraic geometry.
Here, we focus on abelian varieties and closely related varieties (such as Kummer varieties). We can show that the Brauer-Jacobian of an abelian variety $X$ is algebraic whenever it exists, and it exists in very particular circumstances. {restatable*}theoremmain For a complex torus $X$ with $\dim X>1$, the following are equivalent:
  1. (1)
    $X$ is isomorphic to some product $E_{1}\times E_{2}\times\dots\times E_{n}$ for $E_{i}$ pairwise isogenous CM elliptic curves;
  2. (2)
    $X$ is isogenous to $E^{n}$ for a CM elliptic curve $E$;
  3. (3)
    $\mathfrak{C}(X)$ exists as a complex torus;
  4. (4)
    $\mathfrak{C}(X)$ exists as a complex torus and in fact is an abelian variety;
  5. (5)
    $X$ has maximal Picard number.
Moreover, for the Kummer variety associated to an abelian variety we have
{restatable*}
theoremkummercon If $\mathcal{A}$ is a complex abelian variety with an $m$-Jacobian $\mathfrak{C}_{m}(\mathcal{A})$, then the Kummer variety $\mathcal{K}(\mathcal{A})$ associated to $\mathcal{A}$ also has an $m$-Jacobian admitting a canonical isomorphism $\mathfrak{C}_{m}(\mathcal{K}(\mathcal{A}))\cong\mathfrak{C}_{m}(\mathcal{A})$.
We also demonstrate that one can often calculate the isomorphism class of $\mathfrak{C}(X)$. For example, for two elliptic curves $E_{1},E_{2}$ which both have CM by the same order $\mathcal{O}$, we find that
$[\mathfrak{C}(E_{1}\times E_{2})]=[E_{1}][E_{2}]\in\operatorname{Cl}(\mathcal{O}),$
where $[E]$ denotes the element of $\operatorname{Cl}(\mathcal{O})$ corresponding to the isomorphism class of an elliptic curve $E$ with CM by $\mathcal{O}$. If $E_{1},E_{2}$ are isogenous CM elliptic curves, each with CM by an order $\mathcal{O}_{i}$ in the same imaginary quadratic number field $K$, then we show that $\mathfrak{C}(E_{1}\times E_{2})$ is an elliptic curve with CM by the order $\mathcal{O}_{1}+\mathcal{O}_{2}\subseteq K$, the smallest order containing both $\mathcal{O}_{i}$ (Theorem 4.1.4).
More generally, for $X\cong E_{1}\times\dots\times E_{n}$ we have
$\mathfrak{C}(X)\cong\prod_{i<j}\mathfrak{C}(E_{i}\times E_{j}),$
so $\mathfrak{C}(X)$ is a product of $\binom{n}{2}$ pairwise isogenous CM elliptic curves (Theorem 4.1.5).
In fact, for $X$ satisfying the equivalent conditions of Theorem 1.2, not only is $\mathfrak{C}(X)$ an abelian variety, $\mathfrak{C}_{m}(X)$ for all $m$ between $2$ and $\dim X$ is an abelian variety of maximal Picard number (Theorem 4.1.6).
For such varieties $X$, we end up with a complex torus $B$ and a “natural” inclusion $\operatorname{Br}(X)\hookrightarrow B$ as the torsion points. If $B$ is itself algebraic and sufficiently functorial, then one could use this geometric $2$-Jacobian for $\operatorname{Br}(X)$ to study it using the geometric and arithmetic properties of rational points on varieties. This would have many significant consequences, including applications to the Tate conjecture.
As should not be surprising, that is, sadly, not the case. From an abstract point of view, the Hodge structure (and Galois representation) on ${\rm H}^{2}$ is the wrong weight to naturally get an abelian variety. Concretely, we give many examples of abelian varieties $\mathcal{A}$ over fields $k$ whose associated Brauer-Jacobians are not defined over $k$, for example. In fact when $\mathcal{A}=E_{1}\times E_{2}$ is primitive (so the endomorphism rings of the $E_{i}$ and $\mathfrak{C}(\mathcal{A})$ are all equal) we can describe exactly which cases have the same minimal field of definition for $\mathcal{A}$ and $\mathfrak{C}(\mathcal{A})$.
{restatable*}
propositiondecompfieldofdef Assume $\mathcal{A}=\mathbf{C}^{2}/\Lambda$ is a complex torus of dimension $2$ with a Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$ such that $\mathcal{A}$ is primitive. Then there exist elliptic curves $E_{1},E_{2}$ with $\mathcal{A}\simeq E_{1}\times E_{2}$ such that $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))=\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))$ if and only if the group element $[\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O}(\mathcal{A}))$ has order $\leq 2$.
Given that the construction of $\mathfrak{C}_{m}(X)$ appears to be “morally wrong”, it is somewhat bizarre that we keep finding algebraic tori (both here and in [beauville]). Note that there are some cases described in Remark 7.6 of [shioda72] where it is not clear if the torus is algebraic. One is naturally led to the following question.
Question 1.3.
For which values of $m$ is there a proper, smooth, complex variety $X$ admitting an $m$-Jacobian so that $\mathfrak{C}_{m}(X)$ is not algebraic?
In Section 2 we discuss generalities on higher-weight Jacobians for blowups and equivariant $m$-Jacobians, which will be applied to computing $m$-Jacobians for Kummer varieties in Theorem 1.2.
Our computations of the (Brauer-)Jacobian of a complex abelian variety rely on some background from number theory and complex algebraic geometry given in Section 3. More specifically, in Section 3.1 we discuss the connection between complex elliptic curves and lattices in $\mathbf{C}$ in order to define a “lattice product” which computes $\mathfrak{L}_{2}(E_{1}\times E_{2})$, and in Section 3.2 we consider the minimal field of definiton of a complex elliptic curve using class field theory, which will be used for proving Proposition 1.2. With the necessary background material on lattices and complex elliptic curves established, we can define and compute the Brauer-Jacobians for complex abelian varieties by considering them as complex tori $\mathbf{C}^{n}/\Gamma$ in Section 3.3.
We can then consider the structure of higher-weight Jacobians as abelian varieties. We use the class group of the endomorphism algebra of a CM elliptic curve $E$ to compute the $m$-Jacobians for complex abelian varieties isogenous to a power of $E$, beginning with products $E_{1}\times\dots\times E_{n}$ for elliptic curves $E_{i}$ each isogenous to $E$ and $m=2$ in Section 4.1. In Section 4.2 we show that in fact all complex abelian varieties having a Brauer-Jacobian can be decomposed as a product in this way (Theorem 1.2), and furthermore consider how the minimal field of definition for different product decompositions of complex abelian surfaces relates to the $m$-Jacobians. In Section 4.3 we consider when an abelian variety could be isomorphic to its $m$-Jacobian by leveraging a particular decomposition into a product of elliptic curves from [bf68].
Finally in Section 5 we extend our discussion of higher-weight Jacobians for complex abelian varieties to complex K3 surfaces. Our calculations of $j$-invariants for various CM elliptic curves/lattices can be found in Appendix A.
Acknowledgments
We would like to thank Lalita Devadas, János Kollár, and Martin Olsson for helpful conversations and correspondence. Devadas was supported by NSF postdoc grant DMS-2102960. Lieblich was supported by NSF grant DMS-1901933 and NSF FRG grant DMS-2152235.
2. Generalities on higher-weight Jacobians
In this section, we briefly discuss higher-weight Jacobians in general, then discuss applications for blowups and equivariant $m$-Jacobians. Fix a smooth, proper, complex variety $X$.
Lemma 2.1.
Let $V$ be a pure real Hodge structure of weight $m$ with Hodge decomposition $V_{\mathbf{C}}=\bigoplus V^{i,j}$. For any $i$ and $j$ such that $i+j=m$, the $\mathbf{R}$-linear map $V\to V^{i,j}$ (from a real vector space to a complex vector space) is surjective.
Proof.
Write $W^{i,j}=V^{i,j}\oplus V^{j,i}$, and let $W_{0}^{i,j}$ be the set of elements of the form $\alpha+\overline{\alpha}$ for $\alpha\in V^{i,j}$. By basic descent theory, the map $W_{0}^{i,j}\otimes\mathbf{C}\to W^{i,j}$ is an isomorphism. Moreover, the description of the invariants shows that the natural map $W_{0}^{i,j}\to V^{i,j}$ induced by the projection is an isomorphism of real vector spaces. Since $W_{0}^{i,j}\subset V$ (as $V$ is itself identified with the invariants of $V_{\mathbf{C}}$ under conjugation), the result follows. ∎
Corollary 2.2.
Given a pure integral Hodge structure $H$ of weight $m$, the subgroup $\mathfrak{L}(H)\subset H^{0,m}$ is a lattice if and only if one (and hence both) of the following equivalent conditions hold:
  1. (1)
    $\operatorname{rk}\mathfrak{L}(H)=2\dim_{\mathbf{C}}H^{0,m}$;
  2. (2)
    $\dim_{\mathbf{Q}}{\rm im}(H\otimes\mathbf{Q}\to H^{0,m})=2\dim_{\mathbf{C}}H^{0,m}$.
Proof.
By Lemma 2.1, the map $\mathfrak{L}(H)\otimes\mathbf{R}\to H^{0,m}$ is a surjection, which shows that $\mathfrak{L}(H)$ spans the real vector space underlying $H^{0,m}$. Thus, we can choose a saturated sublattice $\Lambda\subset\mathfrak{L}(H)$ that is a real basis for $H^{0,m}$, yielding an identification $\Lambda\otimes\mathbf{R}\to H^{0,m}$. The resulting topological group (torus) $G:=H^{0,m}/\Lambda$ is compact, so any element $\gamma\in\mathfrak{L}(H)\setminus\Lambda$ will have multiples in $G$ that cluster, showing that $\mathfrak{L}(H)$ cannot be discrete in $H^{0,m}$.
The equivalence of the two conditions follows from the fact that the rank of $\mathfrak{L}(H)$ is equal to the dimension of its $\mathbf{Q}$-span inside of $H^{0,m}$. ∎
Thus, there is a purely numerical criterion for $\mathfrak{L}(H)$ to be discrete in $H^{0,m}$. This will be useful for studying the properties of higher-weight Jacobians under natural geometric operations.
Definition 2.3.
Given a pure Hodge structure $H$ of weight $m$, the Jacobian discrepancy of $H$ is the number $\delta(H)=\operatorname{rk}\mathfrak{L}(H)-2\dim_{\mathbf{C}}H^{0,m}$ (see Notation 1.1). If $H={\rm H}^{m}(X,\mathbf{Z})$, we will write $\delta_{m}(X)$ for $\delta(H)$.
Observation 2.4.
We can also read $\delta(H)$ off from the rational Hodge structure $H_{\mathbf{Q}}:=H\otimes_{\mathbf{Z}}\mathbf{Q}$: writing $L(H_{\mathbf{Q}})$ for the image of the map of rational vector spaces $H_{\mathbf{Q}}\to H^{0,m}$, we have $\delta(H)=\dim_{\mathbf{Q}}L(H_{\mathbf{Q}})-2\dim_{\mathbf{C}}H^{0,m}.$
By definition, $H$ has a Jacobian if and only if $\delta(H)=0$. Given a collection of Hodge structures $H_{i}$, we see that $\delta(\bigoplus H_{i})=\sum\delta(H_{i})$.
Lemma 2.5.
Let $H$ be a pure Hodge structure of weight $m$ and let $r=\dim_{\mathbf{C}}H^{0,m}$. For any positive prime number $p$, we have
$\dim_{\mathbf{F}_{p}}\mathfrak{C}(H)[p]=2r+\delta.$
Proof.
Since $\mathfrak{L}(H)$ spans $H^{0,m}$, we can choose a saturated lattice $\Lambda\subset\mathfrak{L}(H)\subset H^{0,m}$. It follows that $\mathfrak{C}(H)$ is isomorphic to $(\mathbf{R}/\mathbf{Z})^{2r}/\Gamma$ for some finitely generated free abelian subgroup $\Gamma$ of non-negative rank equal to $\delta(H)$. By dividing elements of $\Gamma$ by $p$, it follows that we have $\dim_{\mathbf{F}_{p}}\mathfrak{C}(H)[p]=2r+\delta(H)$, as desired. ∎
Corollary 2.6.
If $H_{i}$, $i=1,\ldots,n$, are pure Hodge structures of weight $m$ then $H=\bigoplus H_{i}$ has a Jacobian if and only if each $H_{i}$ has a Jacobian.
Proof.
By construction we have $\mathfrak{C}(H)=\bigoplus\mathfrak{C}(H_{i})$, so the result follows immediately from Lemma 2.5. ∎
Corollary 2.7.
If $P\to X$ is a projective bundle over $X$ of relative dimension $d$, then $P$ has an $m$-Jacobian if and only if $X$ has an $m$-Jacobian.
Proof.
The projective bundle formula gives an isomorphism of Hodge structures
${\rm H}^{m}(P,\mathbf{Z})\cong\bigoplus_{i=0}^{d}{\rm H}^{m-2i}(X,\mathbf{Z})(-i).$
If $H_{i}$ is the weight $m$ Hodge structure ${\rm H}^{m-2i}(X,\mathbf{Z})(-i)$, we note that $H_{i}^{0,m}$ is zero for $i>0$; thus $H_{i}$ has a trivial $m$-Jacobian for $i>0$. The result then follows from Corollary 2.6. ∎
Proposition 2.8.
Suppose $Y\subset X$ is a smooth proper subvariety of codimension $d$ and $\pi:\widetilde{X}\to X$ is the blowup along $Y$. Then $\delta_{m}(\widetilde{X})=\delta_{m}(X)$. Thus, $\widetilde{X}$ has an $m$-Jacobian if and only if $X$ has an $m$-Jacobian.
Moreover, if $\widetilde{X}$ has an $m$-Jacobian then the natural map
$\mathfrak{C}_{m}(X)\to\mathfrak{C}_{m}(\widetilde{X})$
is an isomorphism.
Proof.
Let $E$ be the preimage of $Y$ in $\widetilde{X}$. By Theorem 7.31 of [voisinthing], there is an isomorphism of Hodge structures
${\rm H}^{m}(\widetilde{X},\mathbf{Z})\cong{\rm H}^{m}(X,\mathbf{Z})\oplus\bigoplus_{i=1}^{d-1}{\rm H}^{m-2i}(Y,\mathbf{Z})(-i).$
As in the proof of Corollary 2.7, we note that the weight $m$ Hodge structures $H_{i}:={\rm H}^{m-2i}(Y,\mathbf{Z})(-i)$ have $H_{i}^{0,m}=0$. It follows that $\delta_{m}(\widetilde{X})=\delta_{m}(X)+\sum_{i=1}^{d-1}\delta(H_{i})=\delta_{m}(X)$.
To see the final assertion, note that (since smooth varieties have rational singularities) the natural map ${\rm H}^{m}(X,\mathscr{O}_{X})\to{\rm H}^{m}(\widetilde{X},\mathscr{O}_{\widetilde{X}})$ is an isomorphism that induces a map of $m$-Jacobians $\mathfrak{C}_{m}(X)\to\mathfrak{C}_{m}(\widetilde{X})$. The result follows. ∎
Corollary 2.9.
Given two birational varieties $X$ and $X^{\prime}$ of dimension at most $4$, we have that $X$ has a $q$-Jacobian if and only if $X^{\prime}$ has a $q$-Jacobian, for any $q$.
Proof.
We can compare $X$ and $X^{\prime}$ by birational morphisms having centers of dimension at most $2$. Following the chain of morphisms and using Proposition 2.8 gives the desired result. ∎
Now suppose $H$ is a pure Hodge structure of weight $m$ admitting an action of a finite group $G$ (in the category of Hodge structures).
Lemma 2.10.
If $H$ has a Jacobian then the invariant substructure $H^{G}$ has a Jacobian.
Proof.
By flatness, we have $(H^{G})_{\mathbf{Q}}=(H_{\mathbf{Q}})^{G}$. Thus, it suffices to consider the rational Hodge structure $H_{\mathbf{Q}}$. By semisimplicity, we have a decomposition $H_{\mathbf{Q}}=H_{\mathbf{Q}}^{G}\oplus H^{\prime}$ of Hodge structures. The result now follows from Corollary 2.7. ∎
Proposition 2.11.
Let $X$ be a smooth projective variety with an action by a finite group $G$ such that the quotient $Y=X/G$ is smooth. If $X$ has a $q$-Jacobian then $Y$ has a $q$-Jacobian.
Proof.
By the Cartan-Leray spectral sequence (Theorem XVI.8.4 of [cartaneilenberg]), we have
${\rm H}^{q}(Y,\mathbf{Q})={\rm H}^{q}(X,\mathbf{Q})^{G}.$
The result thus follows from Lemma 2.10. ∎
Corollary 2.12.
Suppose $Y\dashrightarrow X$ is a rational double cover of smooth algebraic surfaces. Then for any $m$, $X$ has an $m$-Jacobian if and only if $Y$ has an $m$-Jacobian.
Proof.
By Corollary 2.9, Proposition 2.11, and the fact that any quadratic field extension (in characteristic different from $2$) is a $\mathbf{Z}/2\mathbf{Z}$-Galois extension, we see that is suffices to show that there is a birational model $\widetilde{X}\to X$ such that the pullback of $Y$ to $\widetilde{X}$ extends to a double cover $\widetilde{Y}\to\widetilde{X}$ with $\widetilde{Y}$ also smooth. To do this, note that we can make a birational modification $X^{\prime}\to X$ so that the normalization of $X^{\prime}$ is the function field of $Y$ gives a finite morphism $Y^{\prime}\to X^{\prime}$ with branch locus $B\subset X^{\prime}$ an snc divisor. If $B$ is smooth then so is $Y^{\prime}$. Otherwise, $B$ has simple crossings of curves, so $Y^{\prime}$ has $A_{1}$-singularities over the singular points of $B$. Blowing up these points in $X^{\prime}$ to yield $\widetilde{X}\to X^{\prime}$, we can explicitly calculate that the normalization $\widetilde{Y}\to\widetilde{X}$ is now smooth. ∎
We end this section with a comment on the relation between higher-weight Jacobians and the framework of [totaro].
Proposition 2.13.
If a pure Hodge structure $H$ of weight $n$ has a Jacobian then there is a decomposition $H_{\mathbf{Q}}=H_{0}\oplus H^{\prime}$ where $H_{0}$ has type $(n,0,\ldots,0,n)$ and $H_{0}^{0,n}=H_{\mathbf{Q}}^{0,n}$.
Proof.
Let $V$ be the $\mathbf{Q}$-vector space given by ${\rm im}(H_{\mathbf{Q}}\to H^{0,n})$. By construction, there is a canonical decomposition $V\otimes\mathbf{C}=H^{n,0}\oplus H^{0,n}$, compatible with complex conjugation and the Hodge symmetry. This defines $H_{0}$ as a quotient of $H_{\mathbf{Q}}$. ∎
3. Complex lattices and complex tori
Lattices in the complex numbers are closely related to complex elliptic curves; this is discussed in various sources including [silverman09]. In particular, by [silverman09, Corollary VI.5.1.1], every complex elliptic curve $E$ is isomorphic as a complex Lie group to $\mathbf{C}/\Lambda$ for some lattice $\Lambda\subset\mathbf{C}$, which is unique up to homothety.
Furthermore, by [silverman09, Theorem VI.4.1], every isogeny of complex elliptic curves $\phi:E_{1}\to E_{2}$ (with $E_{i}=\mathbf{C}/\Lambda_{i}$) arises from a homothety of $\Lambda_{1}$ with a sublattice of $\Lambda_{2}$—i.e., the isogeny corresponds to a nonzero complex number $\alpha$ with $\alpha\Lambda_{1}\subseteq\Lambda_{2}$. We will refer to lattices with this property as isogenous, as we do for the corresponding elliptic curves. Isogeny of lattices is an equivalence relation.
Definition 3.1.
A lattice $\Lambda\subseteq\mathbf{C}$ has complex multiplication or CM if its endomorphism ring ${\rm End}(\Lambda):=\left\{\alpha\in\mathbf{C}:\alpha\Lambda\subseteq\Lambda\right\}$ satisfies $\mathbf{Z}\subsetneq{\rm End}(\Lambda)$.
A discussion of orders in imaginary quadratic number fields and their relation with lattices with complex multiplication can be found in [cox11, Section 7]. It is well-known that the endomorphism ring of a complex lattice with CM is an order $\mathcal{O}$ in an imaginary quadratic number field $K$ [cox11, Lemma 7.2]. Such an order is equal to $\mathbf{Z}+f\mathcal{O}_{K}$ for $f$ the index of $\mathcal{O}$ in $\mathcal{O}_{K}$; this $f$ is also called the conductor of $\mathcal{O}$. It is also well-known that a lattice $\Lambda=\langle 1,\tau\rangle$ has complex multiplication if and only if $\tau$ is the root of a polynomial $ax^{2}+bx+c\in\mathbf{Z}[x]$.
The set of homothety classes of lattices with endomorphism ring $\mathcal{O}$ is in bijection with the ideal class group $\operatorname{Cl}(\mathcal{O})=\mathcal{I}(\mathcal{O})/\mathcal{P}(\mathcal{O})$ (for $\mathcal{I}(\mathcal{O})$ the group of proper/invertible fractional $\mathcal{O}$-ideals under multiplication, and $\mathcal{P}(\mathcal{O})$ the subgroup of principal fractional $\mathcal{O}$-ideals), since a lattice $\Lambda$ with ${\rm End}(\Lambda)=\mathcal{O}$ is homothetic to an ideal of $\mathcal{O}$, and a homothety of (fractional) ideals of $\mathcal{O}$ can be considered as multiplication by a principal fractional ideal.
Lemma 3.2.
If $\mathcal{O}_{1},\mathcal{O}_{2}$ are orders in an imaginary quadratic number field $K$ and $\Lambda_{1},\Lambda_{2}$ are lattices with CM by $\mathcal{O}_{1},\mathcal{O}_{2}$ respectively, then $\Lambda_{1}$ is isogenous to $\Lambda_{2}$.
Proof.
The lattice $\Lambda_{1}$ with CM by $\mathcal{O}_{1}$ is homothetic to an ideal of $\mathcal{O}_{1}$, which is a sublattice of $\mathcal{O}_{1}$. Hence $\Lambda_{1}$ is isogenous to $\mathcal{O}_{1}$ (and similarly for $\Lambda_{2},\mathcal{O}_{2}$). Furthermore, if $\mathcal{O}_{i}$ has conductor $f_{i}$, then both $\mathcal{O}_{1},\mathcal{O}_{2}$ are sublattices of the order in $K$ with conductor $\gcd(f_{1},f_{2})$, so they are also isogenous to one another. It follows that $\Lambda_{1}$ is isogenous to $\Lambda_{2}$. ∎
3.1. Lattice product
The “lattice product” defined below is directly related to the Brauer-Jacobian of the product of two elliptic curves; in Lemma 3.3.2 we will compute that for two isogenous CM elliptic curves $\mathbf{C}/\Lambda_{1},\mathbf{C}/\Lambda_{2}$, the Brauer-Jacobian is isomorphic to the elliptic curve corresponding to this lattice product $\Lambda_{1}\Lambda_{2}$, which we distinguish from the direct product lattice $\Lambda_{1}\times\Lambda_{2}\subseteq\mathbf{C}^{2}$.
Definition 3.1.1.
Given two lattices $\Lambda_{1},\Lambda_{2}\subseteq\mathbf{C}$, the lattice product $\Lambda_{1}\Lambda_{2}$ or $\Lambda_{1}\cdot\Lambda_{2}$ is the additive subgroup of $\mathbf{C}$ consisting of finite sums of products of an element of $\Lambda_{1}$ with an element of $\Lambda_{2}$, or
$\Lambda_{1}\Lambda_{2}=\left\{\sum_{i=1}^{s}a_{i}b_{i}:s\in\mathbf{N},a_{i}\in\Lambda_{1},b_{i}\in\Lambda_{2}\right\}.$
Lemma 3.1.2.
  1. (i)
    The lattice product is commutative and associative.
  2. (ii)
    If $\Lambda_{1}=\langle v_{1},w_{1}\rangle,\Lambda_{2}=\langle v_{2},w_{2}\rangle$, then $\Lambda_{1}\Lambda_{2}=\langle v_{1}v_{2},v_{1}w_{2},w_{1}v_{2},w_{1}w_{2}\rangle$.
  3. (iii)
    If $\Lambda_{1},\Lambda_{3}$ are homothetic, and $\Lambda_{1}\Lambda_{2}$ is a lattice, then $\Lambda_{1}\Lambda_{2},\Lambda_{2}\Lambda_{3}$ are homothetic and the latter is also a lattice.
  4. (iv)
    If $\mathcal{O}={\rm End}(\Lambda)$, then $\mathcal{O}\Lambda=\Lambda$.
  5. (v)
    For two orders $\mathcal{O}_{1},\mathcal{O}_{2}$ of an imaginary quadratic number field $K$ with conductors $f_{1},f_{2}$ respectively, then $\mathcal{O}_{1}\mathcal{O}_{2}=\mathcal{O}_{3}$ for $\mathcal{O}_{3}$ the order of $K$ with conductor $f_{3}=\gcd(f_{1},f_{2})$. ($\mathcal{O}_{3}$ can alternatively be described as the ‘smallest’ order of $K$ that contains both $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$.)
Proof.
Properties (i)-(iii) are all immediate consequences of the definition of the lattice product; properties (iv) and (v) are both simple computations. ∎
Lemma 3.1.3.
If $\Lambda_{1}=\langle v_{1},w_{1}\rangle$ and $\Lambda_{2}=\langle v_{2},w_{2}\rangle$ are isogenous lattices with complex multiplication, then $\Lambda_{1}\Lambda_{2}=\langle v_{1}v_{2},v_{1}w_{2},w_{1}v_{2},w_{1}w_{2}\rangle$ is also a lattice.
Proof.
[…]
Then $\Lambda_{2}$ is homothetic to a sublattice of $\Lambda_{1}$; considering the images of $1$ and $\tau_{2}$ via this homothety, we see that this is equivalent to saying that we have two lattice elements $x_{1}+y_{1}\tau_{1},x_{2}+y_{2}\tau_{1}$ such that their ratio $\frac{x_{1}+y_{1}\tau_{1}}{x_{2}+y_{2}\tau_{1}}=\tau_{2}$.
Since $\Lambda_{1}$ has complex multiplication, $\tau_{1}$ is the root of a quadratic polynomial over $\mathbf{Z}$, so $\mathbf{Q}(\tau_{1})$ is a degree $2$ field extension of $\mathbf{Q}$. Since $\frac{x_{1}+y_{1}\tau_{1}}{x_{2}+y_{2}\tau_{1}}=\tau_{2}$ we see that $\tau_{2}\in\mathbf{Q}(\tau_{1})$. Hence the vector space over $\mathbf{Q}$ generated by $1,\tau_{1},\tau_{2}$ and $\tau_{1}\tau_{2}$ is the field $\mathbf{Q}(\tau_{1})$ and has dimension exactly $2$. As $1,\tau_{1}$ are necessarily linearly independent over $\mathbf{R}$ (hence over $\mathbf{Q}$), we may assume that $\tau_{2},\tau_{1}\tau_{2}$ can be written as $\mathbf{Q}$-linear combinations of $1$ and $\tau_{1}$.
Then we can clear denominators so that for some $a\in\mathbf{Z}$ we have
$\Lambda_{1}=\langle 1,\tau_{1}\rangle\subseteq\langle 1,\tau_{1},\tau_{2},\tau_{1}\tau_{2}\rangle=\Lambda_{1}\Lambda_{2}\subseteq\frac{1}{a}\Lambda_{1}.$
We can deduce that $\Lambda_{1}\Lambda_{2}$ satisfies the necessary discreteness conditions to be a lattice from this ‘sandwich’ $\Lambda_{1}\subseteq\Lambda_{1}\Lambda_{2}\subseteq\frac{1}{a}\Lambda_{1}.$
For ideals and fractional ideals there is also a notion of a product, which would be written the same way as the lattice product; however, there is no ambiguity of notation as we show in the following lemma.
Lemma 3.1.4.
If $\Lambda_{1},\Lambda_{2}$ are fractional ideals of the same order $\mathcal{O}$, then the lattice product $\Lambda_{1}\Lambda_{2}$ is equal to their product as fractional ideals.
Proof.
Both $\Lambda_{1},\Lambda_{2}$ are isogenous to $\mathcal{O}$ as they are homothetic to an ideal (which is a sublattice) of $\mathcal{O}$. Thus they are isogenous CM lattices and their lattice product is a lattice by Lemma 3.1.3.
To see that their lattice product is the same as their product as fractional ideals, it is sufficient to note that any $\mathcal{O}$-linear combination of products $ab$ with $a\in\Lambda_{1},b\in\Lambda_{2}$ is in fact a $\mathbf{Z}$-linear combination of such elements, since $\mathcal{O}$ is contained in the endomorphism rings of both lattices. ∎
Lemma 3.1.5.
Consider two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ of an imaginary quadratic field $K$, and let $\Lambda$ be a proper fractional ideal of $\mathcal{O}$. Then $\mathcal{O}^{\prime}\Lambda$ is a proper fractional ideal of $\mathcal{O}^{\prime}$, and it is isomorphic as an $\mathcal{O}^{\prime}$-module to the extension of scalars $\Lambda\mathbin{\mathop{\otimes}\displaylimits_{\mathcal{O}}}\mathcal{O}^{\prime}$.
Proof.
By Lemma 3.2, $\Lambda_{1}$ is isogenous to $\mathcal{O}$, so we know that $\mathcal{O}^{\prime}\Lambda$ is a lattice by Lemma 3.1.3.
It is also evident that $\mathcal{O}^{\prime}\Lambda$ is closed under multiplication by $\mathcal{O}^{\prime}$ and that it is contained in $K$, so $\mathcal{O}^{\prime}\Lambda$ is a fractional ideal of $\mathcal{O}^{\prime}$. To show that it is proper, we will show that it is invertible.
As $\Lambda$ is a proper fractional ideal of $\mathcal{O}$, it has an inverse $\Lambda^{-1}$—that is, $\Lambda^{-1}$ is another proper (invertible) fractional ideal of $\mathcal{O}$ such that $\Lambda\Lambda^{-1}=\mathcal{O}$, where the product can be taken either as fractional ideals or as lattices by Lemma 3.1.4. (In fact $\Lambda^{-1}=\overline{\Lambda}$.)
Then by Lemma 3.1.2(i), we see that
$(\mathcal{O}^{\prime}\Lambda)(\mathcal{O}^{\prime}\Lambda^{-1})=(\mathcal{O}^{\prime}\mathcal{O}^{\prime})(\Lambda\Lambda^{-1})=\mathcal{O}^{\prime}\mathcal{O}=\mathcal{O}^{\prime},$
and $\mathcal{O}^{\prime}\Lambda^{-1}$ is a fractional ideal of $\mathcal{O}^{\prime}$ for the same reason that $\mathcal{O}^{\prime}\Lambda$ is.
Hence by Lemma 3.1.4, we see that $\mathcal{O}^{\prime}\Lambda$ is an invertible (hence proper) fractional ideal of $\mathcal{O}^{\prime}$ as we have just exhibited an inverse.
There is an obvious map $\Lambda\mathbin{\mathop{\otimes}\displaylimits_{\mathcal{O}}}\mathcal{O}^{\prime}\to\mathcal{O}^{\prime}\Lambda$, given by $x\otimes\alpha\mapsto\alpha x$ for $x\in\Lambda,\alpha\in\mathcal{O}^{\prime}$; it is simple to show that this map is an isomorphism of $\mathcal{O}^{\prime}$-modules. ∎
Lemma 3.1.6.
Let $\Lambda_{1},\Lambda_{2}$ be isogenous lattices with CM, so $\Lambda_{1}\Lambda_{2}$ is a lattice by Lemma 3.1.3. Then
  1. (i)
    $\mathcal{O}_{1}={\rm End}(\Lambda_{1})$ and $\mathcal{O}_{2}={\rm End}(\Lambda_{2})$ are orders in the same imaginary quadratic number field $K$;
  2. (ii)
    ${\rm End}(\Lambda_{1}\Lambda_{2})=\mathcal{O}_{1}\mathcal{O}_{2}=\mathcal{O}_{3}$, another order in $K$.
Proof.
A “converse” to (i) can be found in Lemma 3.2.
Up to homothety, we can assume that for some complex numbers $\tau_{1},\tau_{2}$ we have $\Lambda_{i}=\langle 1,\tau_{i}\rangle$; since $1\in\Lambda_{i}$, we can thus assume that ${\rm End}(\Lambda_{i})\subseteq\Lambda_{i}$ and that $\Lambda_{i}$ is a (proper) fractional ideal of ${\rm End}(\Lambda_{i})$. This is in fact a containment of lattices. Furthermore, in the proof of Lemma 3.1.3, we showed that in this situation $\tau_{2}\in\mathbf{Q}(\tau_{1})$, which is an imaginary quadratic number field because $\Lambda_{1}$ has CM. Then in fact $\mathbf{Q}(\tau_{2})=\mathbf{Q}(\tau_{1})=K$ is then necessarily the fraction field of both ${\rm End}(\Lambda_{i})$. This proves (i).
To prove (ii), we use Lemma 3.1.2(i) and (iv):
$\Lambda_{1}\Lambda_{2}=\Lambda_{1}\mathcal{O}_{1}\cdot\mathcal{O}_{2}\Lambda_{2}=\Lambda_{1}\cdot\mathcal{O}_{3}\cdot\Lambda_{2}=(\mathcal{O}_{3}\Lambda_{1})(\mathcal{O}_{3}\Lambda_{2}).$
By Lemma 3.1.5, both $\mathcal{O}_{3}\Lambda_{i}$ for $i=1,2$ are proper, hence invertible, fractional ideals of $\mathcal{O}_{3}$; applying Lemma 3.1.4, we see that their lattice product is the same as their product as fractional ideals. Therefore $(\mathcal{O}_{3}\Lambda_{1})(\mathcal{O}_{3}\Lambda_{2})$ is also an invertible, hence proper fractional ideal of $\mathcal{O}_{3}$, meaning ${\rm End}(\Lambda_{1}\Lambda_{2})=\mathcal{O}_{3}$ as we desired to show. ∎
3.2. Field of definition
There exists a correspondence between complex elliptic curves and complex lattices, as discussed in previous sections. We note that two elliptic curves defined over number fields are isomorphic over $\mathbf{C}$ if and only if they have the same $j$-invariant; we generally treat the $j$-invariant of an elliptic curve and the $j$-invariant of the associated lattice as interchangeable [silvermanAd, Proposition 4.4].
Definition 3.2.1.
If a complex elliptic curve $E$ is isomorphic to one defined over a number field $L$, then necessarily the $j$-invariant $j(E)\in L$. Conversely, as long as $j\neq 0,1728$, then $E$ is isomorphic (over $\mathbf{C}$) to the curve defined by
$y^{2}=x^{3}+\frac{3j}{1728-j}x+\frac{2j}{1728-j},$
and it is easy to find examples of elliptic curves over $\mathbf{Q}$ with $j$-invariant 00 or $1728$.
Therefore the ‘smallest’ number field over which an elliptic curve isomorphic to $E$ over $\mathbf{C}$ may be defined is $\mathbf{Q}(j(E))$, which we refer to as the field of definition of $E$.
We will give a criterion in Proposition 1.2 for when a complex abelian surface $\mathcal{A}$ having a Brauer-Jacobian $\mathfrak{C}(A)$ can be defined over the same field as the elliptic curve $\mathfrak{C}(A)$.
The first main theorem of complex multiplication [cox11, Theorem 11.1] states that for an order $\mathcal{O}$ of an imaginary quadratic field $K$, the $j$-invariant of any proper fractional $\mathcal{O}$-ideal $I$ (or equivalently, the $j$-invariant of any lattice with CM by $\mathcal{O}$) is an algebraic integer $j(I)$, and for all $I$ the field extension $K(j(I))$ is the ring class field of the order $\mathcal{O}$.
The ring class field $L_{\mathcal{O}}$ has the property that $\operatorname{Cl}(\mathcal{O})\simeq{\rm Gal}(L_{\mathcal{O}}/K)$. Under this group isomorphism, the class of a proper fractional ideal $I$ corresponds to the element $\sigma_{I}$ of the Galois group which permutes $j$-invariants by $[\overline{I}]=[I^{-1}]$—i.e., $\sigma_{I}(j(J))=j(I^{-1}J)=j(\overline{I}J))$ for $I,J$ proper fractional ideals of $\mathcal{O}$ [cox11, Corollary 11.37].
Furthermore, [cox11, Lemma 9.3] states that $L_{\mathcal{O}}$ is Galois over $\mathbf{Q}$ with Galois group the semidirect product $\operatorname{Cl}(\mathcal{O})\rtimes\mathbf{Z}/2\mathbf{Z}$ where $\mathbf{Z}/2\mathbf{Z}$ acts on $\operatorname{Cl}(\mathcal{O})\simeq{\rm Gal}(L_{\mathcal{O}}/K)$ via $\sigma\mapsto\sigma^{-1}$ and $\mathbf{Z}/2\mathbf{Z}$ acts on $L_{\mathcal{O}}$ by complex conjugation.
This description of ${\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$ will allow us to identify when for two elliptic curves $E_{1},E_{2}$ with CM by $\mathcal{O}$ the fields of definition $\mathbf{Q}(j(E_{1}))$ and $\mathbf{Q}(j(E_{2}))$ are equal.
Lemma 3.2.2.
Let $E_{1},E_{2}$ be two nonisomorphic complex elliptic curves with CM by $\mathcal{O}$, an order in an imaginary quadratic field $K$. Let $L_{\mathcal{O}}$ be the ring class field of $\mathcal{O}$, which is a degree $2n$ extension of $\mathbf{Q}$ for $n=|\operatorname{Cl}(\mathcal{O})|$. Then one of the following is true:
  1. (i)
    $\mathbf{Q}(j(E_{1}),j(E_{2}))=L_{\mathcal{O}}$,
  2. (ii)
    or $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))$ is a degree $n$ extension of $\mathbf{Q}$, and in $\operatorname{Cl}(\mathcal{O})$ the element $[E_{1}][E_{2}]^{-1}$ has order $2$.
Proof.
By [cox11, Proposition 13.2], the $j(E_{i})$ have minimal polynomial equal to the class equation $H_{\mathcal{O}}(X)$, which is a degree $n$ polynomial in $\mathbf{Z}[X]$. Hence $n=[\mathbf{Q}(j(E_{i})):\mathbf{Q}]$.
It then follows from the fundamental theorem of Galois theory that the subfields $\mathbf{Q}(j(E_{i}))\subseteq L_{\mathcal{O}}$ are each the fixed field of an order $\frac{2n}{n}=2$ subgroup of ${\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})=\operatorname{Cl}(\mathcal{O})\rtimes\mathbf{Z}/2\mathbf{Z}$—this order $2$ subgroup is ${\rm Gal}(L_{\mathcal{O}}/\mathbf{Q}(j(E_{i})))\subseteq{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$.
Writing $\mathbf{Z}/2\mathbf{Z}$ as {±$\{\pm 1\}$, the order $2$ elements of $\operatorname{Cl}(\mathcal{O})\rtimes\mathbf{Z}/2\mathbf{Z}$ are either $(x,1)$ for $x\in\operatorname{Cl}(\mathcal{O})$ of order $2$, or $([E],-1)$ for any $[E]\in\operatorname{Cl}(\mathcal{O})$. Since $\operatorname{Cl}(\mathcal{O})$ acts transitively and freely on the set of $j$-invariants of elliptic curves with CM by $\mathcal{O}$ [cox11, Corollary 11.37], the subfield $\mathbf{Q}(j(E_{1}))$ cannot arise as the fixed field of $(x,1)$ for any $x\in\operatorname{Cl}(\mathcal{O})$ of order $2$.
Therefore for any elliptic curve $E$ with ${\rm End}(E)=\mathcal{O}$, the subextension $\mathbf{Q}(j(E))\subseteq L_{\mathcal{O}}$ is the fixed field of $([I],-1)$ for some proper fractional ideal $I$ of $\mathcal{O}$. Using the description of the action of $\operatorname{Cl}(\mathcal{O})$ on the $j$-invariants of elliptic curves with CM by $\mathcal{O}$ given in [cox11, Corollary 11.37], we see that for any proper fractional ideal $J$ of $\mathcal{O}$, the element $([I],-1)\in{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$ sends $j(J)$ to $j(\overline{IJ})=j(I^{-1}J^{-1})$. Therefore $([I],-1)$ fixes $j(J)$ if and only if $[J]^{2}=[I]^{-1}\in\operatorname{Cl}(\mathcal{O})$.
Hence $\mathbf{Q}(j(E))$ is the fixed field of $([E]^{-2},-1)\in{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$.
We now consider $\mathbf{Q}(j(E_{1}),j(E_{2}))$, which must lie between the extensions $\mathbf{Q}(j(E_{1}))$ and $L_{\mathcal{O}}$ of $\mathbf{Q}$. By degree reasons,
$[\mathbf{Q}(j(E_{1}),j(E_{2})):\mathbf{Q}]=n\text{ or }2n.$
In the latter case, (i) obviously holds; in the former case, $\mathbf{Q}(j(E_{1}),j(E_{2}))=\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))$, so $([E_{1}]^{-2},-1)=([E_{2}]^{-2},-1)\in{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$. Then $[E_{1}]^{2}=[E_{2}]^{2}\in\operatorname{Cl}(\mathcal{O})$, meaning $[E_{1}][E_{2}]^{-1}$ has order $2$ (since $E_{1}\not\simeq E_{2}$) and we have (ii) in this case, completing the proof. ∎
Example 3.2.3.
[…]
$\mathfrak{O}=\langle 1,3\sqrt{-3}\rangle,\mathcal{L}_{1}=\langle 3,2+\sqrt{-3}\rangle,\mathcal{L}_{2}=\langle 3,1+\sqrt{-3}\rangle$
the $j$-invariants of which generate the distinct degree $3$ extensions
$\mathbf{Q}(j(\mathfrak{O}))$ $=\mathbf{Q}(\sqrt[3]{2}),$
$\mathbf{Q}(j(\mathcal{L}_{1}))$ $=\mathbf{Q}(\zeta_{3}\sqrt[3]{2}),$
$\mathbf{Q}(j(\mathcal{L}_{2}))$ $=\mathbf{Q}(\zeta_{3}^{2}\sqrt[3]{2}),$
as calculated in Example A.3. This is in agreement with Lemma 3.2.2 as $\operatorname{Cl}(\mathfrak{O})$ does not have any elements of order $2$. ∎
Example 3.2.4.
The order $\mathfrak{O}^{\prime}=\mathbf{Z}[\sqrt{-48}]$ in the field $\mathfrak{K}=\mathbf{Q}(\sqrt{-3})$ has ring class field $L_{\mathfrak{O}^{\prime}}=\mathfrak{K}(i,\sqrt{2})=\mathbf{Q}(i,\sqrt{2},\sqrt{3})$ and class group $\operatorname{Cl}(\mathfrak{O}^{\prime})=\mathbf{Z}/2\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. The four homothety classes of lattices with CM by $\mathfrak{O}^{\prime}$ can be computed to be represented by
$\mathfrak{O}^{\prime}=\langle 1,4\sqrt{-3}\rangle,\langle 4,\sqrt{-3}\rangle,\langle 4,2+\sqrt{-3}\rangle,\langle 2,1+2\sqrt{-3}\rangle,$
the $j$-invariants of which each generate the degree $4$ extension $\mathbf{Q}(\sqrt{2},\sqrt{3})/\mathbf{Q}$ (Example A.4). This is in agreement with Lemma 3.2.2 as every element of $\operatorname{Cl}(\mathfrak{O}^{\prime})$ has order $2$. ∎
When we consider abelian varieties that arise as the products of elliptic curves with different endomorphism rings, the following map between class groups will be useful to describe the behavior of the Brauer-Jacobian.
Proposition 3.2.5.
Let $K=\mathbf{Q}(\alpha)$ be an imaginary quadratic field with $\mathcal{O}_{K}=\mathbf{Z}[\alpha]$. For integers $c\mid a$, the map $\phi_{c,a,\alpha}:\operatorname{Cl}(\mathbf{Z}[a\alpha])\to\operatorname{Cl}(\mathbf{Z}[c\alpha])$ defined by $\Lambda\mapsto\mathbf{Z}[c\alpha]\cdot\Lambda$ is a group homomorphism. Furthermore, for $d\mid c\mid a$ we have $\phi_{d,c,\alpha}\circ\phi_{c,a,\alpha}=\phi_{d,a,\alpha}$, and $\phi_{a,a,\alpha}$ is the identity homomorphism.
Proof.
It is clear that $\phi_{c,a,\alpha}$ is well-defined on classes of proper fractional ideals byLemma 3.1.5; in fact it also follows that as modules over $\mathbf{Z}[c\alpha]$ there is an isomorphism $\Lambda\otimes_{\mathbf{Z}[a\alpha]}\mathbf{Z}[c\alpha]\cong\phi_{c,a,\alpha}(\Lambda)$. The rest of the proposition follows from Lemma 3.1.4 and the properties of the lattice product. ∎
In the following lemma we describe the behavior of the $\phi$ map defined in Proposition 3.2.5 with regard to $j$-invariants of class group elements.
Lemma 3.2.6.
Let $K$ be an imaginary quadratic number field with $\mathcal{O}_{K}=\mathbf{Z}[\alpha]$, and let $c\mid a$ be positive integers. Then if $E$ is a complex elliptic curve with CM by $\mathbf{Z}[a\alpha]$, any elliptic curve $E^{\prime}$ corresponding to the class $\phi_{c,a,\alpha}([E])\in\operatorname{Cl}(\mathbf{Z}[c\alpha])$ has $j(E^{\prime})\in\mathbf{Q}(j(E))$.
Proof.
It is not hard to show that the map $\phi_{c,a,\alpha}:\operatorname{Cl}(\mathbf{Z}[a\alpha])\to\operatorname{Cl}(\mathbf{Z}[c\alpha])$ is compatible with the map of Galois groups arising from the inclusion of ring class fields $L_{\mathbf{Z}[c\alpha]}\subseteq L_{\mathbf{Z}[a\alpha]}$. See [cox11, Exercise 9.19] for the inclusion of ring class fields; the compatibility with $\phi$ arises from how the map $\operatorname{Cl}(\mathcal{O})\to{\rm Gal}(L_{\mathcal{O}}/K)$ is defined [cox11, Lemma 5.19, Theorem 11.36].
Let $L_{c}=L_{\mathbf{Z}[c\alpha]},\mathcal{O}_{c}=\mathbf{Z}[c\alpha]$ and define $L_{a},\mathcal{O}_{a}$ similarly. By the proof of Lemma 3.2.2 we see that $\mathbf{Q}(j(E))$ is the fixed field of the order $2$ subgroup of ${\rm Gal}(L_{a}/\mathbf{Q})$ generated by $([E]^{-2},-1)$. The map $\phi_{c,a,\alpha}$ can be extended to a map ${\rm Gal}(L_{a}/\mathbf{Q})\to{\rm Gal}(L_{c}/\mathbf{Q})$, and the image of $([E]^{-2},-1)$ under this map is $([E^{\prime}]^{-2},-1)\in\operatorname{Cl}(\mathcal{O}_{c})\rtimes\mathbf{Z}/2\mathbf{Z}={\rm Gal}(L_{c}/\mathbf{Q})$, which we know fixes $j(E^{\prime})$. Hence $([E]^{-2},-1)$ fixes $j(E^{\prime})$ as well, so $j(E^{\prime})\in\mathbf{Q}(j(E))$ as we desired to show. ∎
Example 3.2.7.
The order $\mathfrak{R}=\mathbf{Z}[\sqrt{-9}]=\mathbf{Z}[3i]$ in $K=\mathbf{Q}(i)$ has ring class field $L_{\mathfrak{R}}=K(\sqrt{3})$ and class group $\operatorname{Cl}(\mathfrak{R})=\mathbf{Z}/2\mathbf{Z}$, and the two homothety classes of lattices with CM by $\mathfrak{R}$ are represented by $\mathfrak{R}=\langle 1,3i\rangle,\langle 3,1+i\rangle$ (Example A.1). For both these lattices, the $j$-invariant generates the quadratic extension $\mathbf{Q}(\sqrt{3})/\mathbf{Q}$.
[…]
The map $\phi_{3,6,i}:\operatorname{Cl}(\mathbf{Z}[6i])\to\operatorname{Cl}(\mathbf{Z}[3i])$ is the surjective map $\mathbf{Z}/4\mathbf{Z}\to\mathbf{Z}/2\mathbf{Z}$; it is easy to check Lemma 3.2.6 in this case is satisfied, as $\mathbf{Q}(\sqrt{3})\subseteq\mathbf{Q}(i\sqrt[4]{12})$ and $\mathbf{Q}(\sqrt{3})\subseteq\mathbf{Q}(\sqrt[4]{12})$. ∎
3.3. Computing cohomology of complex tori
Now that we have the necessary background material on lattices and complex elliptic curves, we can define and compute the Brauer-Jacobians for complex abelian varieties that are a product of isogenous CM elliptic curves. We will compute the cohomology groups of these abelian varieties as complex tori $\mathbf{C}^{n}/\Gamma$.
The following discussion is based on Chapter I of [mumabvar].
[…]
The natural inclusion of the constant sheaf $\underline{\mathbf{Z}}$ in the structure sheaf $\mathscr{O}_{X}$ gives rise to a map of the cohomology groups ${\rm H}^{r}(X,\mathbf{Z})\to{\rm H}^{r}(X,\mathscr{O}_{X})$; it can be shown that this arises as the exterior power of the map
[…]
We will be mostly concerned with the image and cokernel of the map in the case $r=2$.
We now describe the map $\Gamma^{*}\to\overline{T}$ in the one-dimensional case where $\Gamma=\Lambda\subseteq\mathbf{C}$ is a lattice in $\mathbf{C}$ and $X\cong\mathbf{C}/\Lambda$ is a complex elliptic curve.
We can assume that the lattice $\Lambda$ is $\langle 1,\tau\rangle$ for some $\tau$ in the upper half plane (via homothety of the lattice). In this case, $T\simeq\mathbf{C}$, with the map defined by sending $\alpha\in T={\rm Hom}_{\mathbf{C}}(\mathbf{C},\mathbf{C})$ to $\alpha(1)$; we see that $\alpha(x)=x\cdot\alpha(1)$ for all $x\in\mathbf{C}$ by the $\mathbf{C}$-linearity of $\alpha$. Similarly, any $\beta\in\overline{T}={\rm Hom}_{\mathbf{C}-\rm{antilinear}}(\mathbf{C},\mathbf{C})$ satisfies $\beta(x)=\overline{x}\cdot\beta(1)$ for all $x\in\mathbf{C}$, so we also have an isomorphism $\overline{T}\simeq\mathbf{C}$ via $\beta\mapsto\beta(1)$.
Letting $e_{1},e_{2}$ be the $\mathbf{Z}$-basis of $\Lambda^{*}$ which is dual to the $\mathbf{Z}$-basis $\{1,\tau\}$ of $\Lambda$, it is a simple matrix computation to find that
(1) $e_{1}\mapsto\frac{-\tau}{\overline{\tau}-\tau},\;e_{2}\mapsto\frac{1}{\overline{\tau}-\tau}$
via the map $\Lambda^{*}\to T\oplus\overline{T}\twoheadrightarrow\overline{T}\simeq\mathbf{C}$ that we have described.
For a lattice $\Gamma\subseteq\mathbf{C}^{n}$ defining a complex torus $X=\mathbf{C}^{n}/\Gamma$, we see by the above that the cohomology group ${\rm H}^{2}(X,\mathbf{Z})$ is a free abelian group of rank $\binom{2n}{2}$, and ${\rm H}^{2}(X,\mathscr{O})$ is a complex vector space of dimension $\binom{n}{2}$. It is not hard to see that $X$ admits a Brauer-Jacobian precisely when the image of the map ${\rm H}^{2}(X,\mathbf{Z}(1))\to{\rm H}^{2}(X,\mathscr{O})$ has rank $n^{2}-n$.
Notation 3.3.1.
In the above setting, we will write $\mathfrak{L}(\Gamma)$ for $\mathfrak{L}(X)$, and similarly for $\mathfrak{L}_{m}$.
This permits us to think about the space of tori $X$ as parametrized by lattices $\Gamma$ in a fixed copy of $\mathbf{C}^{n}$.
Lemma 3.3.2.
If $X=E_{1}\times E_{2}$ for isogenous CM elliptic curves $E_{i}=\mathbf{C}/\Lambda_{i}$, the lattice $\mathfrak{L}(X)$ is homothetic to the lattice product $\Lambda_{1}\Lambda_{2}$.
Proof.
Up to isomorphism, we can assume that each $\Lambda_{i}=\langle 1,\tau_{i}\rangle$ for $\tau_{i}$ a complex number lying in the upper half plane. Note that by Lemma 3.1.6 both $\Lambda_{1},\Lambda_{2}$ have CM by an order in the imaginary quadratic number field $K=\mathbf{Q}(\tau_{1})=\mathbf{Q}(\tau_{2})$, and $\Lambda_{1}\Lambda_{2}$ is also a lattice with CM by an order in $K$.
Let $\Gamma=\Lambda_{1}\times\Lambda_{2}$ and $\epsilon_{i}=\frac{1}{\overline{\tau_{i}}-\tau_{i}}$. Letting $\{e_{1},e_{2},e_{3},e_{4}\}$ be the basis of $\Gamma^{*}$ dual to the basis $\{\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}\tau_{1}\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right),\left(\begin{smallmatrix}0\\ \tau_{2}\end{smallmatrix}\right)\}$ of $\Gamma$, the map $H^{1}(X,\mathbf{Z})\simeq\Gamma^{*}\simeq\mathbf{Z}^{4}\to\mathbf{C}^{2}\simeq{\rm H}^{1}(X,\mathscr{O})$ is given by
$e_{1}\mapsto\left(\begin{smallmatrix}-\epsilon_{1}\tau_{1}\\ 0\end{smallmatrix}\right),e_{2}\mapsto\left(\begin{smallmatrix}\epsilon_{1}\\ 0\end{smallmatrix}\right),e_{3}\mapsto\left(\begin{smallmatrix}0\\ -\epsilon_{2}\tau_{2}\end{smallmatrix}\right),e_{4}\mapsto\left(\begin{smallmatrix}0\\ \epsilon_{2}\end{smallmatrix}\right)$
using the computation of Equation 1.
We may then calculate the image of the map $\bigwedge^{2}(\mathbf{Z}^{4})\simeq\mathbf{Z}^{6}\to\bigwedge^{2}(\mathbf{C}^{2})\simeq\mathbf{C}$ based on the images of the basis elements $e_{i}\wedge e_{j}$ to be the additive subgroup
$\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})={\epsilon_{1}\epsilon_{2}}\langle 1,\tau_{1},\tau_{2},\tau_{1}\tau_{2}\rangle=\epsilon_{1}\epsilon_{2}\Lambda_{1}\Lambda_{2}\subseteq\mathbf{C},$
which is evidently homothetic to the lattice $\Lambda_{1}\Lambda_{2}$. ∎
Example 3.3.3.
If $E=\mathbf{C}/\langle 1,\tau\rangle$ is an elliptic curve with complex multiplication, it follows that $\mathfrak{C}(E\times E)$ is isomorphic to the elliptic curve $\mathbf{C}/\Lambda^{\prime}$ for $\Lambda^{\prime}=\langle 1,\tau,\tau^{2}\rangle=\langle 1,\tau\rangle\cdot\langle 1,\tau\rangle$. While $\Lambda^{\prime}$ can be shown to be isogenous to $\Lambda$ (hence $\mathfrak{C}(E\times E)$ is isogenous to $E$), it is not usually the case that $\Lambda,\Lambda^{\prime}$ are homothetic, meaning $\mathfrak{C}(E\times E)$ is not isomorphic to $E$.
As an example of this, consider $\Lambda=\langle 1,\frac{1+2i}{3}\rangle$. Then $\Lambda^{\prime}=\langle 1,\frac{1+2i}{3},\frac{-3+4i}{9}\rangle,$ which we can compute to be equal to $\langle\frac{1}{3},\frac{2i}{9}\rangle$, which by visual inspection is not homothetic to $\Lambda$. (We may also compute the $j$-invariants of $\Lambda$ and $\Lambda^{\prime}$ to see that the corresponding elliptic curves are not isomorphic; this is done for the lattices $\langle 3,1+2i\rangle$ and $\langle 3,2i\rangle$, to which $\Lambda$ and $\Lambda^{\prime}$ respectively are homothetic, in Example A.2 of the Appendix.) ∎
Lemma 3.3.4.
If $X=E_{1}\times\dots\times E_{n}$ for isogenous CM elliptic curves $E_{i}=\mathbf{C}/\Lambda_{i}$, the lattice $\mathfrak{L}(X)$ is homothetic to the lattice
$\prod_{1\leq i<j\leq n}\Lambda_{i}\Lambda_{j}\subseteq\mathbf{C}^{\binom{n}{2}}.$
Proof.
As in the proof of Lemma 3.3.2 we can assume $\Lambda_{i}=\langle 1,\tau_{i}\rangle$ for $\tau_{i}$ and set $\Gamma=\Lambda_{1}\times\dots\times\Lambda_{n}$ and $\epsilon_{i}=\frac{1}{\overline{\tau_{i}}-\tau_{i}}$. We also let $\{e_{1},\dots,e_{2n}\}$ be the basis of $\Gamma^{*}$ dual to the specified basis for $\Gamma$. Using the computation of Equation 1 we see that the map $H^{1}(X,\mathbf{Z})\simeq\mathbf{Z}^{2n}\to{\rm H}^{1}(X,\mathscr{O})\simeq\mathbf{C}^{n}$ can be described by
$e_{2i-1}\mapsto-\epsilon_{i}\tau_{i}v_{i},e_{2i}\mapsto\epsilon_{i}v_{i}$
for $\{v_{1},\dots,v_{n}\}$ the standard basis of $\mathbf{C}^{n}$.
The elements $\{v_{i}\wedge v_{j}:1\leq i<j\leq n\}$ form a basis of ${\rm H}^{2}(X,\mathscr{O})\simeq\bigwedge^{2}\left(\mathbf{C}^{n}\right)\simeq\mathbf{C}^{\binom{n}{2}}$. We may then calculate the map ${\rm H}^{2}(X,\mathbf{Z})\simeq\bigwedge^{2}(\mathbf{Z}^{n})=\mathbf{Z}^{\binom{2n}{2}}\to{\rm H}^{2}(X,\mathscr{O})$ on the generators $e_{r}\wedge e_{s}$ of ${\rm H}^{2}(X,\mathbf{Z})$, similarly to Lemma 3.3.2. Thus we can deduce that the image of ${\rm H}^{2}(X,\mathbf{Z})$ is
$\prod_{1\leq i<j\leq n}\epsilon_{i}\epsilon_{j}\Lambda_{i}\Lambda_{j}\subseteq\mathbf{C}^{\binom{n}{2}}.$
B Lemma 3.1.6 each lattice product $\Lambda_{i}\Lambda_{j}$ is a lattice with CM by an order in the field $K=\mathbf{Q}(\tau_{i})$, and the image of ${\rm H}^{2}(X,\mathbf{Z})$ is clearly homothetic to the desired lattice in $\mathbf{C}^{\binom{n}{2}}$. ∎
Lemma 3.3.5.
In the setting of Lemma 3.3.4, the lattice $\mathfrak{L}_{m}(X)$ for $m>2$ is homothetic to the lattice If $X=E_{1}\times\dots\times E_{n}$ for isogenous CM elliptic curves $E_{i}=\mathbf{C}/\Lambda_{i}$, the lattice $\mathfrak{L}(X)$ is homothetic to the lattice
$\prod_{1\leq i_{1}<i_{2}<\dots<i_{m}\leq n}\Lambda_{i_{1}}\Lambda_{i_{2}}\dots\Lambda_{i_{m}}.$
Proof.
We define $\tau_{i},\epsilon_{i},e_{i},v_{i}$ as in the proof of Lemma 3.3.4. The elements $\{v_{i_{1}}\wedge v_{i_{2}}\wedge\dots\wedge v_{i_{m}}:1\leq i_{1}<i_{2}<\dots<i_{m}\leq n\}$ form a basis of ${\rm H}^{m}(X,\mathscr{O})\simeq\bigwedge^{m}\left(\Gamma^{*}\right)\simeq\mathbf{C}^{\binom{n}{m}}$. We may then calculate the map ${\rm H}^{m}(X,\mathbf{Z})\simeq\bigwedge^{m}(\mathbf{Z}^{n})=\mathbf{Z}^{\binom{2n}{m}}\to{\rm H}^{m}(X,\mathscr{O})$ on the generators $e_{j_{1}}\wedge e_{j_{2}}\wedge\dots\wedge e_{j_{m}}$ of ${\rm H}^{m}(X,\mathbf{Z})$ to be the additive subgroup
$\prod_{1\leq i_{1}<i_{2}<\dots<i_{m}\leq n}\epsilon_{i_{1}}\epsilon_{i_{2}}\dots\epsilon_{i_{m}}\Lambda_{i_{1}}\Lambda_{i_{2}}\dots\Lambda_{i_{m}}\subseteq\mathbf{C}^{\binom{n}{m}}.$
We can apply Lemma 3.1.6 repeatedly to see that each lattice product $\Lambda_{i_{1}}\Lambda_{i_{2}}\dots\Lambda_{i_{m}}$ is a lattice in $\mathbf{C}$ with CM by an order in $K=\mathbf{Q}(\tau_{i})$, and the image of ${\rm H}^{2}(X,\mathbf{Z})$ is clearly homothetic to the desired lattice in $\mathbf{C}^{\binom{n}{m}}$. ∎
4. Structure of Brauer-Jacobians and complex abelian varieties
In this section, we study the structure of Brauer-Jacobians for complex abelian varieties. As we show in Section 4.1, a product of elliptic curves that are all isogenous to a fixed curve $E$ has a Brauer-Jacobian that can be calculated explicitly using the class group of the endomorphism algebra of $E$. Then, in Section 4.2, we show that the abelian varieties with Brauer-Jacobians are precisely those that have such a decomposition. The result is a complete description of $\mathfrak{C}(X)$ for any abelian variety of maximal Picard number.
4.1. Computing the Brauer-Jacobian with class groups
By Lemma 3.3.2 we know that $\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})$ is homothetic to $\Lambda_{1}\Lambda_{2}$ for isogenous CM lattices $\Lambda_{1},\Lambda_{2}$. We can say something further in the case where our pair of isogenous CM lattices have the same endomorphism ring. This ring $R$ must necessarily be an order in an imaginary quadratic number field; we write $\operatorname{Cl}(R)$ for its ideal class group.
Definition 4.1.1.
There is a bijective correspondence between the set of isomorphism classes of elliptic curves with CM by $R$ and $\operatorname{Cl}(R)$. Namely, if our elliptic curve $E$ is $\mathbf{C}/\Lambda$ for the lattice $\Lambda=\langle 1,\tau\rangle$ and ${\rm End}(E)={\rm End}(\Lambda)=R$, then $\Lambda\subseteq{\rm Frac}(R)$ must be an invertible (proper) fractional ideal of $R$. We will write $[E]$ or $[\Lambda]$ for the element of $\operatorname{Cl}(R)$ associated to $E=\mathbf{C}/\Lambda$ via this correspondence.
Lemma 4.1.2.
For two CM elliptic curves $E_{1},E_{2}$ with ${\rm End}(E_{1})={\rm End}(E_{2})=R$, the ring $R$ is an order in an imaginary quadratic number field $K$ with class group $\operatorname{Cl}(R)$, and $E_{i}=\mathbf{C}/\Lambda_{i}$ with $\Lambda_{i}$ a proper fractional ideal of $R$. Then $\mathfrak{C}(E_{1}\times E_{2})$ is a complex elliptic curve such that
$[E_{1}][E_{2}]=[\mathfrak{C}(E_{1}\times E_{2})]\in\operatorname{Cl}(R).$
Proof.
By Lemma 3.3.2 we know that $\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})$ is homothetic to $\Lambda_{1}\Lambda_{2}$. By Lemma 3.1.4 it is clear that $[\Lambda_{1}][\Lambda_{2}]=[\Lambda_{1}\Lambda_{2}]=[\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})]\in\operatorname{Cl}(R)$, from which it follows that $[\mathfrak{C}(E_{1}\times E_{2})]=[E_{1}][E_{2}]\in\operatorname{Cl}(R)$. ∎
In the specific case of $E_{1}=E_{2}$, it follows from Lemma 4.1.2 that
$[\mathfrak{C}(E\times E)]=[E]^{2}\in\operatorname{Cl}({\rm End}(E)).$
We can use this to give an example of an elliptic curve $E$ for which $\mathfrak{C}(E\times E)$ has a different field of definition than $E$.
Example 4.1.3.
We recall from Example 3.2.7 that the $j$-invariant of the lattice $\langle 3,2i\rangle$ (which has endomorphism ring $\mathbf{Z}[6i]$) is real and generates the degree $4$ extension $\mathbf{Q}(\sqrt[4]{12})$ of $\mathbf{Q}$; however, the $j$-invariant of $\langle 3,1+2i\rangle$ (with the same endomorphism ring) is not real and generates the extension $\mathbf{Q}(i\sqrt[4]{12})/\mathbf{Q}$. Furthermore, $\mathbf{Q}(i\sqrt[4]{12})\cap\mathbf{R}=\mathbf{Q}(\sqrt{3})$ and $j(\langle 3,2i\rangle)\notin\mathbf{Q}(\sqrt{3})$, so $j(\langle 3,2i\rangle)\notin\mathbf{Q}(j(\langle 3,1+2i\rangle))$.
If $E$ is the elliptic curve $\mathbf{C}/\langle 3,1+2i\rangle\cong\mathbf{C}/\langle 1,\frac{1+2i}{3}\rangle$ discussed in Example 3.3.3, we see that while $E$ is $\mathbf{C}$-isomorphic to an elliptic curve defined over $\mathbf{Q}(i\sqrt[4]{12})$, the same is not true of the elliptic curve $\mathfrak{C}(E\times E)\simeq_{\mathbf{C}}\mathbf{C}/\langle 3,2i\rangle$, since $j(\mathfrak{C}(E\times E))\notin\mathbf{Q}(i\sqrt[4]{12})$. ∎
For two CM elliptic curves $E_{1},E_{2}$ which are isogenous, we can deduce from the corresponding case for lattices, Lemma 3.1.6(i), that ${\rm End}(E_{1})$ and ${\rm End}(E_{2})$ are orders in the same imaginary quadratic field $K=\mathbf{Q}(\alpha)$. Assuming that $\alpha$ is chosen so $\mathcal{O}_{K}=\mathbf{Z}[\alpha]$, then for some integers $a,b$ we have ${\rm End}(E_{1})=\mathbf{Z}[a\alpha],{\rm End}(E_{2})=\mathbf{Z}[b\alpha]$. We will use the map $\phi$ defined in Proposition 3.2.5 in our general version of Lemma 4.1.2.
For two isogenous CM elliptic curves $E_{1},E_{2}$, there exists an imaginary quadratic number field $K$ and integers $a$ and $b$ such that $\mathcal{O}_{K}=\mathbf{Z}[\alpha]$, ${\rm End}(E_{1})=\mathbf{Z}[a\alpha]$, and ${\rm End}(E_{2})=\mathbf{Z}[b\alpha]$. Let $c=\gcd(a,b)$.
Theorem 4.1.4.
With the immediately preceding notation, $\mathfrak{C}(E_{1}\times E_{2})$ is a complex elliptic curve and we have
$\phi_{c,a,\alpha}([E_{1}])\phi_{c,b,\alpha}([E_{2}])=[\mathfrak{C}(E_{1}\times E_{2})]\in\operatorname{Cl}(\mathbf{Z}[c\alpha]).$
Thus, in the particular case where ${\rm End}(E_{1})={\rm End}(E_{2})$ (so $a=b=c$), we have $[E_{1}][E_{2}]=[\mathfrak{C}(E_{1}\times E_{2})]\in\operatorname{Cl}(\mathbf{Z}[a\alpha]).$
Proof.
As before, we assume $E_{i}=\mathbf{C}/\Lambda_{i}$ with $\Lambda_{i}$ a proper fractional ideal of ${\rm End}(E_{i})$. Via Lemma 3.1.6(ii) and the homothety of $\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})$ and $\Lambda_{1}\Lambda_{2}$ we see that ${\rm End}(\mathfrak{C}(E_{1}\times E_{2}))=\mathbf{Z}[c\alpha]$ (where $c={\rm gcd}(a,b)$ as in the statement of the theorem).
Since $\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})$ is homothetic to $\Lambda_{1}\Lambda_{2}$, it suffices to show that $\Lambda_{1}\Lambda_{2}=(\mathbf{Z}[c\alpha]\cdot\Lambda_{1})(\mathbf{Z}[c\alpha]\cdot\Lambda_{2})$ by the definition of $\phi$, which is computed in the proof of Lemma 3.1.6. ∎
Hence the case where $E_{1},E_{2}$ have different endomorphism rings can be reduced to the case where $a\mid b$—i.e., where ${\rm End}(E_{1})\supseteq{\rm End}(E_{2})$—by considering the maps $\phi_{a,b,\alpha}$.
This allows us to describe $\mathfrak{C}(X)$ and $\mathfrak{C}_{m}(X)$ for $X$ a product of $n$ pairwise isogenous CM elliptic curves.
Theorem 4.1.5.
Let $X\cong E_{1}\times\dots\times E_{n}$, where $E_{i}\cong\mathbf{C}/\Lambda_{i}$ are pairwise isogenous CM elliptic curves. Then
$\mathfrak{C}(X)=\prod_{1\leq i<j\leq n}\mathfrak{C}(E_{i}\times E_{j}).$
Proof.
We can deduce from Lemmas 3.3.2 and 3.3.4 that
$\mathfrak{L}(\Lambda_{1}\times\dots\times\Lambda_{n})=\prod_{1\leq i<j\leq n}\mathfrak{L}(\Lambda_{i}\times\Lambda_{j})\subseteq\mathbf{C}^{\binom{n}{2}};$
the theorem follows. ∎
Theorem 4.1.6.
For $X$ as in Theorem 4.1.5 and $2\leq m\leq n$, we have
$\mathfrak{C}_{m}(X)\cong\prod_{1\leq i_{1}<i_{2}<\dots<i_{m}\leq n}\mathfrak{C}_{m}(E_{i_{1}}\times E_{i_{2}}\times\dots\times E_{i_{m}}).$
In the case $m=n$, if ${\rm End}(E_{i})=\mathcal{O}_{i}=\mathbf{Z}[f_{i}\alpha]$ is the order of conductor $f_{i}$ in the imaginary quadratic number field $K=\mathbf{Q}\otimes\mathcal{O}_{i}=\mathbf{Q}(\alpha)$ for each $i$, then we have
$\left[\mathfrak{C}_{n}(E_{1}\times\dots\times E_{n})\right]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(\mathcal{O}_{0})$
for $\mathcal{O}_{0}=\mathbf{Z}[d\alpha]$ the order in $K$ of conductor $d=\gcd(f_{1},\dots,f_{n})$.
Proof.
As in Lemma 3.3.5, we can compute
$\mathfrak{C}_{m}(X)$ $\cong\prod_{1\leq i_{1}<i_{2}<\dots<i_{m}\leq n}\mathbf{C}/\left(\Lambda_{i_{1}}\Lambda_{i_{2}}\dots\Lambda_{i_{m}}\right)$
$\cong\prod_{1\leq i_{1}<i_{2}<\dots<i_{m}\leq n}\mathfrak{C}_{m}(E_{i_{1}}\times E_{i_{2}}\times\dots\times E_{i_{m}}).$
If $m=n$, then $\mathfrak{C}_{n}(E_{1}\times\dots\times E_{n})\cong\mathbf{C}/(\Lambda_{1}\dots\Lambda_{n})$, and the lattice product
$[\Lambda_{1}\dots\Lambda_{n}]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(\mathcal{O}_{0})$
by repeatedly applying Theorem 4.1.4. ∎
Remark 4.1.7.
In the case $m=n$, we see that $\mathfrak{C}_{m}(E_{1}\times\dots\times E_{m})\cong\mathbf{C}/(\Lambda_{1}\dots\Lambda_{m})$, so in fact
$\mathfrak{C}_{m}(E_{1}\times\dots\times E_{m})$ $\cong\mathfrak{C}(\mathfrak{C}_{m-1}(E_{1}\times\dots\times E_{m-1})\times E_{m})$
$\cong\mathfrak{C}(\mathfrak{C}(\mathfrak{C}_{m-2}(E_{1}\times\dots\times E_{m-2})\times E_{m-1})\times E_{m})$
[…]
$\cong\underbrace{\mathfrak{C}(\mathfrak{C}(\dots\mathfrak{C}(\mathfrak{C}(}_{m-1\text{ times}}E_{1}\times E_{2})\times E_{3})\times\dots)\times E_{m}).$
4.2. Which complex tori have Brauer-Jacobians?
We have been able to describe the Brauer-Jacobian of a product of pairwise isogenous CM elliptic curves in the previous sections. It has been shown ([sm74] for the case of surfaces, [schoen] for the general case) that a complex abelian variety which is isogenous to a product of isogenous CM elliptic curves must be isomorphic to such a product. In the following theorem we use this fact to show that among complex tori, only such products of elliptic curves have Brauer-Jacobians.
[…]
Proof.
It is immediately clear that (1) implies (2); the reverse implication is given by [schoen, Satz 2.4]. It is also immediate that (4) implies (3).
Assuming (1), we may apply Theorem 4.1.5 to see that $\mathfrak{C}(X)$ exists and is a product of complex elliptic curves, hence an abelian variety. Thus (1) implies (4).
Now assume (3). The fact that $\mathfrak{C}(X)={\rm H}^{2}(X,\mathscr{O})/\mathfrak{L}(X)$ is a complex torus implies that $\mathfrak{L}(X)$, the image of ${\rm H}^{2}(X,\mathbf{Z})$ in ${\rm H}^{2}(X,\mathscr{O})\simeq\mathbf{C}^{\binom{n}{2}}$, is a free abelian group of rank $n^{2}-n$. Since ${\rm H}^{2}(X,\mathbf{Z})$ is a free abelian group of rank $\binom{2n}{2}=2n^{2}-n$, this means that the kernel of ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ is isomorphic to $\mathbf{Z}^{n^{2}}$.
By [mumabvar, Chapter I.2, Theorem of Appell-Humbert], this kernel is isomorphic to the Neron-Severi group ${\rm NS}(X)\simeq\rm{Pic}(X)/\rm{Pic}^{0}(X)$, which is always a free abelian group of finite rank (see either [blabvar, Corollary 5.3.9] or [mumabvar, Chapter IV.19, Corollary 2, Theorem 3]). Furthermore, the rank of the Neron-Severi group is $n^{2}$ exactly in the situation of (2) [blabvar, Proposition 5.2.1, Corollary 5.3.8]. Therefore (3) implies (2).
Then (2) implies (1) implies (4) implies (3) implies (2). By [blabvar, Proposition 5.2.1, Corollary 5.3.8], (5) is equivalent to (2). Thus, all conditions are equivalent, as desired. ∎
Remark 4.2.1.
We note that for any complex variety $X$, the cokernel of the map ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ is a complex torus when the Neron-Severi group ${\rm NS}(X)$ is of ‘maximal’ rank; in [beauville] this is referred to as $X$ being $\rho$-maximal. In particular a complex abelian variety is $\rho$-maximal if and only if it satisfies the equivalent conditions of Theorem 1.2.
In fact for complex abelian varieties which are $\rho$-maximal and hence of the form $X\cong E_{1}\times\dots\times E_{n}$ for pairwise isogenous CM elliptic curves $E_{i}\cong\mathbf{C}/\Lambda_{i}$, not only is the cokernel of the map ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ a complex torus, the same is true of the map ${\rm H}^{m}(X,\mathbf{Z})\to{\rm H}^{m}(X,\mathscr{O})$ by Lemma 3.3.5. (This is because each of the lattice products $\Lambda_{i_{1}}\dots\Lambda_{i_{m}}$ is also a lattice in this situation.) We can compute $\mathfrak{C}_{m}(X)$ for $\rho$-maximal $X$ and $2\leq m\leq\dim(X)$ by Theorem 4.1.6.
Definition 4.2.2.
Similarly to the term $\rho$-maximal of [beauville], we say a complex variety $X$ is $q$-maximal if the cokernel of the map ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ is a complex torus, so $X$ has a $q$-Jacobian. We see that $X$ is $2$-maximal if and only if it is $\rho$-maximal if and only if it has a Brauer-Jacobian.
The proof in [schoen, Satz 2.4] of the fact that any complex abelian variety which is isogenous to $E^{n}$ for a CM elliptic curve $E$ is in fact isomorphic to a product of elliptic curves is ‘algebraic’, as it is a consequence of [bass, Theorem 1.7], which states that for a certain class of rings (to which all orders in imaginary quadratic number fields belong) any torsion-free module can be decomposed into a direct sum of rank $1$ modules. However, this decomposition is not unique; it follows that for $X\simeq E_{1}\times\dots\times E_{n}$, the $E_{i}$ are also not unique. This ambiguity of decomposition is discussed in [sm74] for the case $n=2$ using binary quadratic forms, which we restate here. .
Theorem 4.2.3.
If $\mathcal{A}$ is a $2$-maximal complex abelian surface with $\mathcal{A}=\mathbf{C}^{2}/\Gamma$, then $\mathcal{A}\simeq\mathbf{C}/\mathcal{O}(\mathcal{A})\times\mathfrak{C}(\mathcal{A})$ for for $\mathcal{O}(\mathcal{A})=\{z\in\mathbf{C}:z\Gamma\subseteq\Gamma\}$. It then follows that
  1. (i)
    $\mathcal{A}\simeq\mathcal{A}^{\prime}$ for two $2$-maximal complex abelian surfaces if and only if $\mathcal{O}(\mathcal{A})=\mathcal{O}(\mathcal{A}^{\prime})$ and $\mathfrak{C}(\mathcal{A})\simeq\mathfrak{C}(\mathcal{A}^{\prime})$;
  2. (ii)
    and for $E_{i}=\mathbf{C}/\Lambda_{i},i=1,\dots,4$ pairwise isogenous CM elliptic curves with CM by the imaginary quadratic field $K$, we have $E_{1}\times E_{2}\simeq E_{3}\times E_{4}$ if and only if ${\rm End}(E_{1})\cap{\rm End}(E_{2})={\rm End}(E_{3})\cap{\rm End}(E_{4})$ and $\Lambda_{1}\Lambda_{2}=\Lambda_{3}\Lambda_{4}$.
Proof.
Condition (ii) is proved in [sm74, Proposition 4.5].
If $\mathcal{A}=\mathbf{C}^{2}/\Gamma$ has a Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$, by Theorem 1.2 we may assume without loss of generality that $\mathcal{A}=\mathbf{C}/\Lambda_{1}\times\mathbf{C}/\Lambda_{2}$ for $\Lambda_{1},\Lambda_{2}$ lattices in $\mathbf{C}$ with CM by orders $\mathcal{O}_{1},\mathcal{O}_{2}$ in an imaginary quadratic number field $K$ (as $\Lambda_{1},\Lambda_{2}$ are isogenous). It is then clear that $\mathcal{O}(\mathcal{A})=\mathcal{O}_{1}\cap\mathcal{O}_{2}$, and $\mathcal{O}(\mathcal{A})$ is the order in $K$ with conductor $F={\rm lcm}(f_{1},f_{2})$ for $f_{i}$ the conductors of $\mathcal{O}_{i}$. Also set $f_{0}=\gcd(f_{1},f_{2})$ (and note that $F=f_{1}f_{2}/f_{0}$).
Due to [schoen] (and similarly to [bf68, Lemma 8]) we see that the lattices $\Lambda_{1}\times\Lambda_{2}\subseteq\mathbf{C}^{2}$ and $\mathcal{O}(\mathcal{A})\times\Lambda_{1}\Lambda_{2}$ are isomorphic to one another as $\mathcal{O}(\mathcal{A})$-modules, which proves the main statement of the corollary, from which (i) follows.
Remark 4.2.4.
Using the notation of Theorem 4.2.3, we see that the degree of primitivity $m=F/f_{0}$ of the binary quadratic form associated by Shioda and Mitani to $\mathcal{A}$ is also an invariant of $\mathcal{A}$, and in fact if ${\rm End}(\mathfrak{C}(\mathcal{A}))=\mathbf{Z}[\alpha]$ we have $\mathcal{O}(\mathcal{A})=\mathbf{Z}[m\alpha]$. Thus a $2$-maximal complex abelian surface $\mathcal{A}$ is classified up to isomorphism by $\mathfrak{C}(A)$ and $m$. We will refer to $\mathcal{A}$ such that $m=1$ as primitive.
The uniqueness of the decomposition in Theorem 4.2.3 can also be derived from the general case [bf68, Theorem 8], though in this case where $n=2$ we can derive it more directly as we have done here. The general case proved in [bf68] will be useful when we consider the case $n=3$ in Section 4.3.
We can also use Theorem 4.2.3 to apply the results of Section 3.2 concerning the ‘field of definition’ of elliptic curves to the question of when a product $\mathcal{A}=E_{1}\times E_{2}$ of elliptic curves with CM by an order $\mathcal{O}$ and its Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$ can both be defined over a field ‘smaller’ than the ring class field $L_{\mathcal{O}}$ (in the primitive case).
[…]
Proof.
It follows from Theorem 4.2.3(ii) that the condition that ${\rm End}(\mathfrak{C}(\mathcal{A}))=\mathcal{O}(\mathcal{A})$ for $\mathcal{A}$ a $2$-maximal complex abelian surface is equivalent to the condition that for any $E_{1},E_{2}$ with $\mathcal{A}\simeq E_{1}\times E_{2}$ we have ${\rm End}(E_{1})={\rm End}(E_{2})=\mathcal{O}(\mathcal{A})$. We set $\mathcal{O}=\mathcal{O}(\mathcal{A})$ for ease of writing in this proof. In fact, for any $E_{1},E_{2}$ with CM by $\mathcal{O}$, we have $\mathcal{A}\simeq E_{1}\times E_{2}$ if and only if $[E_{1}][E_{2}]=[\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O})$ by Theorem 4.2.3 and Theorem 4.1.4. We can then apply Lemma 3.2.2 to see that $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))=\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))$ if and only if $[E_{1}][E_{2}]^{-1},[E_{1}],$ and $[E_{2}]$ all have order $\leq 2$. Since $\operatorname{Cl}(\mathcal{O})$ is abelian, this implies that $[\mathfrak{C}(\mathcal{A})]=[E_{1}][E_{2}]$ has order $\leq 2$ also. This completes one direction of the proof.
If $[\mathfrak{C}(\mathcal{A})]$ has order $\leq 2$, then $\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))=\mathbf{Q}(j(\mathbf{C}/\mathcal{O}))$, so we can set $E_{1}=\mathfrak{C}(\mathcal{A}),E_{2}=\mathbf{C}/\mathcal{O}$ to get the desired decomposition.∎
Remark 4.2.5.
Note that in Proposition 1.2, the group element $[\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O})$ has order $\leq 2$ if and only if $j(\mathfrak{C}(\mathcal{A}))$ is real. (This is true since one can show from the definition of the $j$-invariant of a lattice that $j(L)=\overline{j(\overline{L})}$, and if $L$ is a proper fractional ideal of an order $\mathcal{O}$, we have $[L][\overline{L}]=[\mathcal{O}]=1\in\operatorname{Cl}(\mathcal{O})$. See [cox11, Theorem 10.9, Exercise 11.1].)
Example 4.2.6.
Recall from Example 4.1.3 that the order $\mathfrak{R}^{\prime}=\mathbf{Z}[\sqrt{-36}]$ in the field $K=\mathbf{Q}(i)$ has ring class field $L_{\mathfrak{R}}^{\prime}=K(\sqrt[4]{12})$ and class group $\operatorname{Cl}(\mathfrak{R}^{\prime})=\mathbf{Z}/4\mathbf{Z}$, with the identity being $1=[\mathfrak{R}^{\prime}]$, and generator $[\langle 3,1+2i\rangle]$; we can compute $[\langle 3,1+2i\rangle]^{2}=[\langle 3,2i\rangle]$.
Let $\mathcal{A}$ be the abelian surface $\mathcal{A}=(\mathbf{C}/\langle 3,1+2i\rangle)^{2}$. Then $\mathfrak{C}(\mathcal{A})\simeq\mathbf{C}/\langle 3,2i\rangle$; note that $j(\langle 3,2i\rangle)\notin\mathbf{Q}(j(\langle 3,1+2i\rangle))=\mathbf{Q}(i\sqrt[4]{12})$ (see Example A.2). It hence appears that the minimal field of definition of $\mathcal{A}$ is distinct from the minimal field of definition of $\mathfrak{C}(\mathcal{A})$.
However, by Theorem 4.2.3(ii), we see that $\mathcal{A}\cong\mathbf{C}/\mathfrak{R}^{\prime}\times\mathbf{C}/\langle 3,2i\rangle$, and $\mathbf{Q}(j(\mathfrak{R}^{\prime}))=\mathbf{Q}(j(\langle 3,2i\rangle))=\mathbf{Q}[\sqrt[4]{12}]$. Therefore $\mathcal{A}$ can be defined over the minimal field of definition of $\mathfrak{C}(\mathcal{A})$, as we would expect from Proposition 1.2 since the order of $[\mathfrak{C}(\mathcal{A})]=[\langle 3,2i\rangle]$ is $2$. ∎
Example 4.2.7.
Recall from Example 3.2.3 that the order $\mathfrak{O}=\mathbf{Z}[\sqrt{-27}]$ in the field $\mathbf{Q}(\zeta_{3})$ has ring class field $L_{\mathfrak{O}}=\mathbf{Q}(\zeta_{3},\sqrt[3]{2})$ and class group $\operatorname{Cl}(\mathfrak{O})=\mathbf{Z}/3\mathbf{Z}$, with the identity being $1=[\mathfrak{O}]$, and generator $[\Lambda_{1}]=[\langle 3,2+\sqrt{-3}\rangle]$; we can compute $[\langle 3,2+\sqrt{-3}\rangle]^{2}=[\langle 3,1+\sqrt{-3}\rangle]=[\Lambda_{2}]$.
Let $\mathcal{A}^{\prime}$ be the abelian surface $\mathcal{A}^{\prime}=(\mathbf{C}/\Lambda_{1})^{2}$. Then $\mathfrak{C}(\mathcal{A}^{\prime})\simeq\mathbf{C}/\langle 3,1+\sqrt{-3}\rangle$; similarly to the previous example, note that $j(\Lambda_{2})\notin\mathbf{Q}(j(\Lambda_{1}))=\mathbf{Q}(\zeta_{3}\sqrt[3]{2})$ (see Example A.3). It hence appears that the minimal field of definition of $\mathcal{A}^{\prime}$ is distinct from the minimal field of definition of $\mathfrak{C}(\mathcal{A}^{\prime})$.
By Theorem 4.2.3(ii), we see that the only other way to decompose $\mathcal{A}^{\prime}$ as a product of two elliptic curves is as $\mathcal{A}^{\prime}=(\mathbf{C}/\Lambda_{1})^{2}\cong\mathbf{C}/\mathfrak{O}\times\mathbf{C}/\Lambda_{2},$ but $j(\mathfrak{O})\notin=\mathbf{Q}(j(\Lambda_{2}))=\mathbf{Q}(\zeta_{3}^{2}\sqrt[3]{2})$—as we would expect from Proposition 1.2, since the order of $[\mathfrak{C}(\mathcal{A}^{\prime})]=[\Lambda_{2}]$ is $3$, we can’t define $\mathcal{A}^{\prime}$ over the minimal field of definition of $\mathfrak{C}(\mathcal{A}^{\prime})$. ∎
4.3. When is an abelian variety isomorphic to its $m$-Jacobian?
If $X$ is a $2$-maximal complex abelian variety, then by Theorem 1.2 we can assume $X=E_{1}\times\dots\times E_{n}$ is a product of $n$ elliptic curves which are CM and pairwise isogenous; by Theorem 4.1.6 we see that $X$ is $m$-maximal for all $m\leq\dim X$ and $\mathfrak{C}_{m}(X)$ is a product of $\binom{n}{m}$ pairwise isogenous CM elliptic curves. It is evident that we can only have $X\simeq\mathfrak{C}_{m}(X)$ for $m=n-1$.
Before considering for which $X=E_{1}\times E_{2}\times\dots\times E_{n}$ we have $X\simeq\mathfrak{C}_{n-1}(X)$, we will wish to check which invariants of $X$ classify it up to isomorphism. Similar to Theorem 4.2.3, where we showed that a $2$-maximal complex abelian surface $\mathcal{A}$ is uniquely classified up to isomorphism by $\mathfrak{C}(A)$ and the degree of primitivity $m$ (Remark 4.2.4) these invariants will be $\mathfrak{C}_{n}(X)$ and $n-1$ integers.
Lemma 4.3.1.
Let $X=E_{1}\times E_{2}\times\dots\times E_{n}$ be a product of CM elliptic curves with each endomorphism ring ${\rm End}(E_{i})$ equal to an order $\mathcal{O}_{i}$ in a fixed imaginary quadratic number field $K$ (hence the elliptic curves are pairwise isogenous). If $f_{i}$ is the conductor of $\mathcal{O}_{i}$ in $\mathcal{O}_{K}=\mathbf{Z}[\alpha]$, let $d=\gcd(f_{1},\dots,f_{n})$ and $N={\rm lcm}(f_{1},\dots,f_{n})$. Then there exist unique positive integers $r_{2},\dots,r_{n-1}$ with $d=r_{1}\mid r_{2}\mid r_{3}\mid\dots\mid r_{n-1}\mid r_{n}=N$ such that if $R_{i}$ is the order in $K$ of conductor $r_{i}$ we have an isomorphism
$X\cong\mathbf{C}/R_{n}\times\mathbf{C}/R_{n-1}\times\dots\times\mathbf{C}/R_{2}\times\mathfrak{C}_{n}(X).$
Proof.
We have already determined $R_{1},R_{n}$ since $r_{1}=d,r_{n}=N$. It remains to determine the $r_{i}$ for $12\leq i\leq n-$ 1.
We note that $[\mathfrak{C}_{n}(X)]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$ by Theorem 4.1.6. Thus by repeatedly applying Theorem 4.2.3 to $X$, we can show that
$X\cong\mathbf{C}/R_{n}\times\mathbf{C}/R_{n-1}\times\dots\times\mathbf{C}/R_{2}\times\mathfrak{C}_{n}(X)$
for orders $R_{2},\dots,R_{n-1}$ such that $R_{n}\subseteq R_{n-1}\subseteq\dots\subseteq R_{2}\subseteq R_{1}$, and we can take $r_{i}$ to be the conductor of $R_{i}$. The uniqueness of this decomposition is a consequence of [bf68, Theorem 8]. Namely, given the endomorphism algebra $K$, the lattice $\Gamma$ and hence the complex torus $X=\mathbf{C}^{n}/\Gamma$ are determined up to isomorphism by the sequence of conductors $d=r_{1},r_{2},\dots,r_{n}=N$ and $\mathfrak{C}_{n}(X)$ (up to isomorphism/homothety of $\mathfrak{C}_{n}(X)$). The lemma follows. ∎
Remark 4.3.2.
In the case $n=2$, we see that $r_{1}=d,r_{2}=N=md$.
We can now describe when $X\simeq\mathfrak{C}(X)$, or the case $n=3$.
Proposition 4.3.3.
For $X$ a $2$-maximal complex abelian threefold, we have $X\simeq\mathfrak{C}(X)$ if and only if $X\cong(\mathbf{C}/\mathcal{O}_{1})^{3}$.
Proof.
Since $X$ is a $2$-maximal complex abelian threefold, by Theorem 1.2 we have $X\cong E_{1}\times E_{2}\times E_{3}$ a product of three elliptic curves with ${\rm End}(E_{i})=\mathcal{O}_{i}$ for orders $\mathcal{O}_{i}$ in an imaginary quadratic number field $K$ (hence the elliptic curves have CM and are pairwise isogenous)
Let $K,d,N,f_{i}$ be as in Lemma 4.3.1. If $r=f_{1}f_{2}f_{3}/(dN)$ we can see easily that $d\mid r\mid N$ and that
$X=E_{1}\times E_{2}\times E_{3}\simeq\mathbf{C}/R_{3}\times\mathbf{C}/R_{2}\times\mathfrak{C}_{3}(X),$
where $R_{1},R_{2},R_{3}$ are the orders in $\mathcal{O}_{K}$ with conductors $d,r=f_{1}f_{2}f_{3}/(dN),N$ respectively, and $\mathfrak{C}_{3}(X)$ is an elliptic curve with CM by $R_{1}$ such that $[\mathfrak{C}_{3}(X)]=\prod_{i=1}^{3}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$.
Applying Theorem 4.1.5 we see that
$\mathfrak{C}(X)\simeq\mathbf{C}/R_{2}\times\mathfrak{C}_{3}(X)\times\mathfrak{C}_{3}(X)\simeq\mathbf{C}/R_{2}\times\mathbf{C}/R_{1}\times E,$
for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{3}(X)]^{2}\in\operatorname{Cl}(R_{1})$.
By Lemma 4.3.1, we see that $X\simeq\mathfrak{C}(X)$ if and only if $E\simeq\mathfrak{C}_{3}(X),N=r,$ and $r=d$. The latter two equalities imply that $f_{1}=f_{2}=f_{3}$. This also means that all the orders $R_{i},\mathcal{O}_{i}$ are all equal to each other.
It then follows that $X\simeq\mathfrak{C}(X)$ if and only if $f_{1}=f_{2}=f_{3}$ and $[E]=[\mathfrak{C}_{3}(X)]\in\operatorname{Cl}(R_{1})=\operatorname{Cl}(\mathcal{O}_{1})$; since then $[\mathfrak{C}_{3}(X)]=[E_{1}][E_{2}][E_{3}]$ and $[E]=[\mathfrak{C}_{3}(X)]^{2}$, it follows that $[\mathfrak{C}_{3}(X)]$ is the identity in $\operatorname{Cl}(\mathcal{O}_{1})$, so $\mathfrak{C}_{3}(x)\cong\mathbf{C}/\mathcal{O}_{1}$. Because in this case, $X\cong\mathbf{C}/\mathcal{O}_{1}\times\mathbf{C}/\mathcal{O}_{1}\times\mathfrak{C}_{3}(X)$, we see that $X\simeq\mathfrak{C}(X)$ is equivalent to $X\cong(\mathbf{C}/\mathcal{O}_{1})^{3}$ as we desired to show. ∎
Remark 4.3.4.
By Proposition 4.3.3 and Lemma 4.3.1 we see that $X=E_{1}\times E_{2}\times E_{3}\cong\mathfrak{C}(X)$ if and only if ${\rm End}(E_{1})={\rm End}(E_{2})={\rm End}(E_{3})$ and $\mathfrak{C}_{3}(X)\cong\mathbf{C}/({\rm End}(E_{1}))$. Such $X$ are isomorphic to $\mathfrak{C}_{3}(X)^{3}$. In other words, we see that a complex abelian threefold $X$ is isomorphic to $\mathfrak{C}(X)$ if and only if $X\cong E^{3}$ for a complex elliptic curve $E$ which is isomorphic to $\mathbf{C}/\mathcal{O}$ for an order $\mathcal{O}$ in an imaginary quadratic number field $K$.
It therefore makes sense to ask which elliptic curves $E$ satisfy this condition, or, in other words, which lattices $\Lambda$ in $\mathbf{C}$ are homothetic to an order $\mathcal{O}$. In order for a lattice $\Lambda=\langle 1,\tau\rangle$ to be an order, it must be a ring, for which it is sufficient to have $\tau^{2}\in\Lambda$—in other words, $\tau^{2}+m\tau+n=0$ for some integers $m,n$, so $\tau$ is an algebraic integer of degree $2$. In fact this is a necessary and sufficient condition for $\langle 1,\tau\rangle$ to be an order in the imaginary quadratic number field $K=\mathbf{Q}(\tau)$.
Definition 4.3.5.
The $\mathfrak{C}_{2}=\mathfrak{C}$ operator takes $2$-maximal complex abelian threefolds to $2$-maximal complex abelian threefolds; thus we can say that $\mathfrak{C}$ acts on the set of isomorphism classes of $2$-maximal complex abelian threefolds.
Remark 4.3.6.
In Proposition 4.3.3 and Remark 4.3.4 we characterize the fixed points of the $\mathfrak{C}$ action.
Proposition 4.3.7.
The orbits of the $\mathfrak{C}$ action on $2$-maximal complex abelian threefolds are finite.
Proof.
Consider some $2$-maximal complex abelian threefold $X\cong E_{1}\times E_{2}\times E_{3}$, with $K,d,N,f_{i}$ defined as in Lemma 4.3.1. Furthermore set $r,R_{i}$ as in the proof of Proposition 4.3.3.
It’s computed in the proof of Proposition 4.3.3 that $X\cong\mathbf{C}/R_{3}\times\mathbf{C}/R_{2}\times\mathfrak{C}_{3}(X)$ and $\mathfrak{C}(X)\simeq\mathbf{C}/R_{2}\times\mathbf{C}/R_{1}\times E$ for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{3}(X)]^{2}\in\operatorname{Cl}(R_{1})$. A very similar computation using Theorem 4.1.5 and Theorem 4.2.3 shows that
$\mathfrak{C}(\mathfrak{C}(X))\cong\mathbf{C}/R_{1}\times E\times E\cong(\mathbf{C}/R_{1})^{2}\times E^{\prime}$
for $E^{\prime}$ an elliptic curve with CM by $R_{1}$ such that $[E^{\prime}]=[E]^{2}=[\mathfrak{C}_{3}(X)]^{4}\in\operatorname{Cl}(R_{1})$.
In fact, for $k\geq 2$ we see that applying $\mathfrak{C}$ to $X$ a total of $k$ times is isomorphic to $(\mathbf{C}/R_{1})^{2}\times E_{(k)}$ where $E_{(k)}$ is an an elliptic curve with CM by $R_{1}$ such that
$[E_{(k)}]=[E]^{\left(2^{k-1}\right)}=[\mathfrak{C}_{3}(X)]^{\left(2^{k}\right)}\in\operatorname{Cl}(R_{1}).$
Since $\operatorname{Cl}(R_{1})$ is finite, up to isomorphism there are only finitely many isomorphism classes $[E_{(k)}]$. Thus there are finitely many isomorphism classes of abelian threefolds of maximal Picard rank in the orbit of $X$ for the $\mathfrak{C}$ action. ∎
Proposition 4.3.8.
For a $2$-maximal complex abelian variety $X$, the ($n-1$)-Jacobian $\mathfrak{C}_{n-1}(X)$ is isomorphic to $X$ if and only if $X\cong(\mathbf{C}/\mathcal{O}_{1})^{n-1}\times E$ for an elliptic curve $E$ with CM by $\mathcal{O}_{1}$ and $[E]\in\operatorname{Cl}(\mathcal{O}_{1})$ having order dividing $n-2$.
Proof.
Since $X$ is $2$-maximal we have $X=E_{1}\times E_{2}\times\dots\times E_{n}$ a product of CM elliptic curves with ${\rm End}(E_{i})=\mathcal{O}_{i}$ for orders $\mathcal{O}_{i}$ in an imaginary quadratic number field $K$. Let $K,d,N,f_{i}$ be as in Lemma 4.3.1; by this Lemma we obtain unique integers $r_{1},\dots,r_{n}$ with $r_{1}=d\mid r_{2}\mid\dots\mid r_{n-1}\mid r_{n}=N$ and an isomorphism
$X=E_{1}\times\dots\times E_{n}\simeq\mathbf{C}/R_{n}\times\dots\times\mathbf{C}/R_{2}\times\mathfrak{C}_{n}(X)$
where $R_{i}$ is the order in $K$ with conductor $r_{i}$, and $\mathfrak{C}_{n}(X)$ is an elliptic curve with CM by $R_{1}$ such that $[\mathfrak{C}_{n}(X)]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$ by Theorem 4.1.6.
Applying Theorem 4.1.6 we see that
$\mathfrak{C}_{n-1}(X)\simeq\mathbf{C}/R_{2}\times\left(\mathfrak{C}_{n}(X)\right)^{n-1}\simeq\mathbf{C}/R_{2}\times(\mathbf{C}/R_{1})^{n-2}\times E,$
for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{n}(X)]^{n-1}\in\operatorname{Cl}(R_{1})$.
By Lemma 4.3.1, we see that $X\simeq\mathfrak{C}_{n-1}(X)$ if and only if $E\simeq\mathfrak{C}_{n}(X),N=r_{2},$ and $d=r_{1}=r_{2}=\dots=r_{n-1}$. The latter two equalities imply that $N=d$, so $f_{1}=f_{2}=\dots=f_{n}$. This also means that all the orders $R_{i},\mathcal{O}_{i}$ are all equal to each other.
It then follows that $X\simeq\mathfrak{C}_{n-1}(X)$ if and only if $f_{1}=\dots=f_{n}$ and $[E]=[\mathfrak{C}_{n}(X)]\in\operatorname{Cl}(R_{1})=\operatorname{Cl}(\mathcal{O}_{1})$; since $[E]=[\mathfrak{C}_{n}(X)]^{n-1}$, it follows that $[\mathfrak{C}_{n}(X)]$ has order dividing $n-2$ in $\operatorname{Cl}(\mathcal{O}_{1})$. Because in this case, $X\cong\left(\mathbf{C}/\mathcal{O}_{1}\right)^{n-1}\times\mathfrak{C}_{n}(X)$, we see that $X\simeq\mathfrak{C}_{n-1}(X)$ is equivalent to $X\cong(\mathbf{C}/\mathcal{O}_{1})^{n-1}\times E$ for an elliptic curve $E$ with CM by $\mathcal{O}_{1}$ and $[E]\in\operatorname{Cl}(\mathcal{O}_{1})$ having order dividing $n-2$, as we desired to show. ∎
Remark 4.3.9.
By Proposition 4.3.8 and Lemma 4.3.1 we see that $X=E_{1}\times\dots\times E_{n}\cong\mathfrak{C}_{n-1}(X)$ if and only if ${\rm End}(\mathfrak{C}_{n}(X))={\rm End}(E_{1})=\dots={\rm End}(E_{n})$ and $[\mathfrak{C}_{n}(X)]^{n-2}\in\operatorname{Cl}({\rm End}(\mathfrak{C}_{n}(X)))$ is the identity.
Definition 4.3.10.
Similarly to the $\mathfrak{C}$ action, we note that the $\mathfrak{C}_{n-1}$ operator takes $2$-maximal complex abelian $n$-folds to $2$-maximal complex abelian $n$-folds; thus we can say that $\mathfrak{C}_{n-1}$ acts on the set of isomorphism classes of $2$-maximal complex abelian $n$-folds.
Proposition 4.3.11.
The orbits of the $\mathfrak{C}$ action on $2$-maximal complex abelian $n$-folds are finite.
Proof.
Consider some $2$-maximal complex abelian $n$-fold $X\cong E_{1}\times E_{2}\times E_{3}$, with $K,d,N,f_{i}$ defined as in Lemma 4.3.1. Furthermore set $r_{i},R_{i}$ as in the proof of Proposition 4.3.3.
It’s computed in the proof of Proposition 4.3.8 that $X\cong\mathbf{C}/R_{n}\times\dots\times\mathbf{C}/R_{2}\times\mathfrak{C}_{n}(X)$ and $\mathfrak{C}_{n-1}(X)\simeq\mathbf{C}/R_{2}\times\left(\mathfrak{C}_{n}(X)\right)^{n-1}\simeq\mathbf{C}/R_{2}\times(\mathbf{C}/R_{1})^{n-2}\times E,$ for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{n}(X)]^{n-1}\in\operatorname{Cl}(R_{1})$.
A very similar computation shows that
$\mathfrak{C}_{n-1}(\mathfrak{C}_{n-1}(X))\cong\mathbf{C}/R_{1}\times E^{n-1}\cong(\mathbf{C}/R_{1})^{n-1}\times E^{\prime}$
for $E^{\prime}$ an elliptic curve with CM by $R_{1}$ such that $[E^{\prime}]=[E]^{n-1}=[\mathfrak{C}_{n-1}(X)]^{\left((n-1)^{2}\right)}\in\operatorname{Cl}(R_{1})$.
In fact, for $k\geq 2$ we see that applying $\mathfrak{C}{n-1}$ to $X$ a total of $k$ times is isomorphic to $(\mathbf{C}/R_{1})^{n-1}\times E_{(k)}$ where $E_{(k)}$ is an an elliptic curve with CM by $R_{1}$ such that
$[E_{(k)}]=[E]^{\left((n-1)^{k-1}\right)}=[\mathfrak{C}_{3}(X)]^{\left((n-1)^{k}\right)}\in\operatorname{Cl}(R_{1}).$
Since $\operatorname{Cl}(R_{1})$ is finite, up to isomorphism there are only finitely many isomorphism classes $[E_{(k)}]$. Thus there are finitely many isomorphism classes of abelian $n$-folds of maximal Picard rank in the orbit of $X$ for the $\mathfrak{C}_{n-1}$ action. ∎
5. Kummer varieties and singular K3 surfaces
In this section, we discuss higher-weight Jacobians for Kummer varieties and singular K3 surfaces, which are related to Kummer varieties of abelian varieties with maximal Picard number.
Definition 5.1.
If $\mathcal{A}$ is an abelian variety, The Kummer variety of an abelian variety $\mathcal{A}$, denoted $\operatorname{\mathcal{K}}(\mathcal{A})$, is the minimal resolution $\mathcal{K}(\mathcal{A})\to\mathcal{A}/\iota$ of the quotient by the involution $\iota:\mathcal{A}\to\mathcal{A},v\mapsto-v$.
Remark 5.2.
The Kummer variety can alternatively be described as the quotient of the blowup at the $2$-torsion points $\widetilde{\mathcal{A}}\to\mathcal{A}$ by a lift $\tilde{\iota}$ of the involution (cf. [huybrechts, Example 3.3]). If $\mathcal{A}$ is an abelian surface, the Kummer surface $\mathcal{K}(\mathcal{A})$ is a K3 surface.
[…]
Proof.
By Proposition 2.8, the blow up $\widetilde{\mathcal{A}}$ of $\mathcal{A}$ at the $2$-torsion points has an $m$-Jacobian. By Proposition 2.11 and Remark 5.2, so does $\operatorname{\mathcal{K}}(\mathcal{A})$. (The case of Brauer-Jacobians for Kummer surfaces was also shown in the proof of [si77, Theorem 6].) ∎
Corollary 5.3.
A singular complex K3 surface has a Brauer-Jacobian.
Proof.
Let $X$ be a singular K3 surface, that is, a K3 surface with Picard number $20$. By [si77], $X$ is rationally a double cover of $\operatorname{\mathcal{K}}(E\times E^{\prime})$ for two isogenous CM elliptic curves $E$ and $E^{\prime}$. Applying Corollary 2.12 gives the result. ∎
Remark 5.4.
In [si77], Shioda and Inose define an elliptic curve $C_{X}$ associated to a singular K3 or abelian surface $X$ which coincides with our definition of the Brauer-Jacobian $\mathfrak{C}(X)$. Furthermore, they proved (via a correspondence with equivalence classes of binary quadratic forms) that any singular complex K3 surface $Y$ (of maximal Picard rank 20) is a double cover of a Kummer surface $X=\mathcal{K}(\mathcal{A})$ associated to an abelian surface $\mathcal{A}$ of maximal Picard rank, and that $\mathfrak{C}(Y)\cong\mathfrak{C}(X)$ [si77, Theorem 4].
Remark 5.5.
Using the discussion of [si77], we see that any singular K3 surface $Y$ which is a double cover of the Kummer surface $\mathcal{K}(\mathcal{A})$ can be defined over the extension of the ring class field $L_{\mathcal{O}(\mathcal{A})}$ by the coordinates of the 2-torsion points of the elliptic curves $\mathbf{C}/\mathcal{O}(\mathcal{A}),\mathfrak{C}(\mathcal{A})\simeq\mathfrak{C}(Y)$, since $\mathcal{A}\cong\mathbf{C}/\mathcal{O}(\mathcal{A})\times\mathfrak{C}(Y)$ (Theorem 4.2.3) and $j(\mathfrak{C}(Y))\in L_{\mathcal{O}(\mathcal{A})}$ (Lemma 3.2.6). The field of definition of a singular K3 surface is discussed in [k3fieldofdef].
Appendix A Calculating $j$-invariants
To compute the class group of an order and the homothety classes of lattices with CM by that order, we used the bijection between the form class group and the ideal class group described in [cox11, Theorem 7.7].
Since Sage has implemented composition of quadratic forms and has methods to list all primitive positive definite forms of a given discriminant, the structure of the form class group can be easily computed. Then by the bijection between the form class group and the ideal class group described in [cox11, Theorem 7.7] (namely, a primitive positive definite quadratic form $f(x,y)=ax^{2}+bxy+cy^{2}$ of discriminant $D$ corresponds to the proper ideal $(a,(-b+\sqrt{D})/2)$ of the unique order $\mathcal{O}_{D}$ with discriminant $D$) we can then compute the homothety classes of lattices with CM by $\mathcal{O}_{D}$.
Sage has also implemented a database of Hilbert class polynomials (the class equation mentioned in the proof of Lemma 3.2.2). If $H_{\mathcal{O}}(X)$ is the Hilbert class polynomial of an order $\mathcal{O}$, we see by [cox11, Proposition 13.2] that by adjoining one root of $H_{\mathcal{O}}(X)$ to the fraction field of $\mathcal{O}$, we obtain the ring class field of $\mathcal{O}$. By then using Sage again to factor various polynomials over this ring class field, we can obtain a simpler description of $L_{\mathcal{O}}$.
Once we have a simple description of $L_{\mathcal{O}}$, we can factor $H_{\mathcal{O}}(X)$ over $L_{\mathcal{O}}$ to get exact values of the $j$-invariants of the lattices with CM by $\mathcal{O}$. To identify which $j$-invariant corresponds to which lattice, we need only then approximate the $j$-invariants of each lattice; this can be done by truncating the $q$-series described in [wa08, Proposition 9.12].
Thus we may compute the following, using $\sim$ to denote homothety of lattices.
Example A.1.
The order $\mathbf{Z}[3i]=\mathbf{Z}[\sqrt{-9}]$ in $\mathbf{Q}(i)$ has discriminant $-36$; the class group is isomorphic to $\mathbf{Z}/2\mathbf{Z}$, with the two classes being represented by the quadratic forms $x^{2}+9y^{2},2x^{2}+2xy+5y^{2}$ or the lattices $\langle 1,3i\rangle,\langle 2,-1+3i\rangle\sim\langle 3,1+i\rangle$, with $j$-invariants
$j(\langle 1,3i\rangle)$ $=76771008+44330496\sqrt{3},$
$j(\langle 3,1+i\rangle)$ $=76771008-44330496\sqrt{3}.$
The ring class field is $L_{\mathbf{Z}[3i]}=\mathbf{Q}(i,\sqrt{3})$. ∎
Example A.2.
The order $\mathbf{Z}[6i]=\mathbf{Z}[\sqrt{-36}]$ in $\mathbf{Q}(i)$ has discriminant $-144$; the class group is isomorphic to $\mathbf{Z}/4\mathbf{Z}$, with the four classes being represented by the quadratic forms (and corresponding lattices)
$x^{2}+36y^{2}:$ $\langle 1,6i\rangle,$
$5x^{2}\pm 4xy+8y^{2}:$ $\langle 5,\pm 2+6i\rangle\sim\langle 3,1\pm 2i\rangle,$
$4x^{2}+9y^{2}:$ $\langle 4,6i\rangle\sim\langle 3,2i\rangle,$
or the lattices  with $j$-invariants
[…]
The ring class field is $L_{\mathbf{Z}[6i]}=\mathbf{Q}(i,\sqrt[4]{12})$. ∎
Example A.3.
The order $\mathbf{Z}[3\sqrt{-3}]=\mathbf{Z}[\sqrt{-27}]$ in $\mathbf{Q}(\sqrt{-3})$ has discriminant $-27$; the class group is isomorphic to $\mathbf{Z}/3\mathbf{Z}$, with the three classes being represented by the quadratic forms (and corresponding lattices)
$x^{2}+27y^{2}:$ $\langle 1,3\sqrt{-3}\rangle,$
$4x^{2}+2xy+7y^{2}$ $:\langle 4,-1+3\sqrt{-3}\rangle\sim\langle 3,2+\sqrt{-3}\rangle,$
$4x^{2}-2xy+7y^{2}$ $:\langle 4,1+3\sqrt{-3}\rangle\sim\langle 3,1+\sqrt{-3}\rangle$
with $j$-invariants
[…]
where $\zeta_{3}=\frac{-1+\sqrt{-3}}{2}$ is a cube root of unity.
The ring class field is $L_{\mathbf{Z}[3\sqrt{-3}]}=\mathbf{Q}(\sqrt{-3},\sqrt[3]{2})$. ∎
Example A.4.
The order $\mathbf{Z}[4\sqrt{-3}]=\mathbf{Z}[\sqrt{-48}]$ in $\mathbf{Q}(\sqrt{-3})$ has discriminant $-48$; the class group is isomorphic to $\mathbf{Z}/2\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$, with the four classes being represented by the quadratic forms (and corresponding lattices)
$x^{2}+48y^{2}:$ $\langle 1,4\sqrt{-3}\rangle,$
$7x^{2}+2xy+7y^{2}:$ $\langle 7,-1+4\sqrt{-3}\rangle\sim\langle 4,2+\sqrt{-3}\rangle,$
$4x^{2}+4xy+13y^{2}:$ $\langle 4,-2+4\sqrt{-3}\rangle\sim\langle 2,1+2\sqrt{-3}\rangle,$
$3x^{2}+16y^{2}:$ $\langle 3,4\sqrt{-3}\rangle\sim\langle 4,\sqrt{-3}\rangle,$
with $j$-invariants
[…]
The ring class field is $L_{\mathbf{Z}[4\sqrt{-3}]}=\mathbf{Q}(i,\sqrt{2},\sqrt{3})$. ∎
[…]